### v1

# Towards Explainable Bilingual Multimodal Misinformation Detection and Localization

Abstract
The increasing realism of multimodal content has made misinformation more subtle and more complex to detect, especially in news media, where images are frequently paired with bilingual subtitles (e.g., Chinese-English). Such content often includes localized image edits and cross-lingual inconsistencies that jointly distort meaning while remaining superficially plausible. We introduce BiMi, a bilingual multimodal framework that jointly performs region-level localization, cross-modal and cross-lingual consistency detection, and natural language explanation for misinformation analysis. To support generalization, BiMi integrates an online retrieval module that supplements model reasoning with up-to-date external context. We further release BiMiBench, a large-scale and comprehensive benchmark constructed by systematically editing real news images and subtitles, comprising 104,000 samples with realistic manipulations across visual and linguistic modalities. To enhance interpretability, we apply Group Relative Policy Optimization (GRPO) to improve explanation quality, marking the first use of GRPO in this domain. Extensive experiments demonstrate that BiMi outperforms strong baselines by up to +8.9 in classification accuracy, +15.9 in localization accuracy, and +2.5 in explanation BERTScore, advancing state-of-the-art performance in realistic, multilingual misinformation detection. Code, models, and datasets are avaliable at https://github.com/RaffeLegend/misinformation.
1 Introduction
The rapid advancement of generative models [1, 2, 3, 4, 5] has drastically lowered the barrier to producing highly realistic multimodal content [6]. While enabling creative applications, this progress has also escalated the proliferation of misinformation that intricately blends visual and textual elements [7, 8]. A particularly insidious form emerges in news media, where images are frequently overlaid with bilingual subtitles (e.g., Chinese-English). Here, subtle visual tampering and cross-lingual inconsistencies are exploited to distort facts while maintaining surface-level plausibility. This becomes especially dangerous for monolingual audiences, who may only comprehend one language and thus remain oblivious to manipulations in the other. A stark example occurred during early 2020, when a CDC report confirming the first U.S. case of COVID-19 community transmission [^1] https://edition.cnn.com/2020/02/27/health/us-cases-coronavirus-community-transmission/index.html was mistranslated by some Chinese social media as asserting that the virus originated in the United States. This misleading bilingual manipulation fueled significant public misunderstanding and geopolitical tensions. Such cases illustrate how multimodal misinformation increasingly leverages localized image edits and asymmetric subtitle translations to manipulate public perception across language communities. Addressing such sophisticated attacks demands more than conventional binary misinformation detection. It requires systems capable of region-level localization of manipulations, identification of cross-modal and cross-lingual inconsistencies, and generation of coherent, cross-lingual natural language explanations—capabilities that remain underexplored.
Despite encouraging progress [7, 8, 9, 10, 11], existing systems exhibit three critical limitations. First, most rely on global image-text matching, lacking the fine-grained region-level localization necessary to expose subtle manipulations. Second, their explanations are often superficial and generic, providing little actionable insight into why content is misleading. Third, they neglect the modeling of bilingual subtitle inconsistencies, overlooking the semantic divergence arising from intentionally misleading translations. These shortcomings underline the need for a unified framework that reasons at a fine granularity across visual and bilingual textual modalities.
Figure 1: Comparison of traditional vs. bilingual misinformation detection tasks. Traditional tasks focus on visual-text consistency with limited outputs (left). Our setting uses tampered news images and bilingual subtitles, enabling richer outputs including region localization, cross-modal consistency, and explanation (right). Red indicates error, green indicates correctness. Best viewed in color.
However, building such a framework is technically challenging. Region localization demands precise visual-textual alignment that is beyond the capacity of most pretrained MLLMs. Generating faithful explanations in bilingual, multimodal settings requires accurate identification and articulation of inconsistencies across modalities and languages. Moreover, cross-lingual reasoning introduces unique complexities, as even fluent translations may introduce subtle but impactful semantic shifts that obscure the misinformation signal. These challenges are further exacerbated in the news domain, characterized by rapidly evolving, diverse, and context-dependent content.
To tackle these challenges, we introduce BiMi, the first bilingual multimodal misinformation detection framework that seamlessly integrates region-level manipulation localization, cross-modal and cross-lingual consistency detection, and natural language explanation generation. An overview of our task and outputs is illustrated in Figure 1. To enhance generalization to emerging events, BiMi is equipped with an online retrieval module that augments model reasoning with real-time external knowledge. Furthermore, to facilitate systematic evaluation, we construct BiMiBench, the first large-scale benchmark for this setting, comprising 104,000 samples with realistic manipulations of news images and bilingual subtitles. Through extensive experiments, BiMi achieves state-of-the-art performance, surpassing strong baselines by +8.9 in classification accuracy, +15.9 in localization accuracy, and +2.5 in explanation BERTScore.
Our contributions are summarized as follows:
  • We release BiMiBench, the first large-scale benchmark for bilingual multimodal misinformation detection, containing 104K samples with fine-grained image manipulations and bilingual subtitle inconsistencies.
  • We propose BiMi, a unified framework that detects image-subtitle misinformation via region localization, cross-modal consistency detection, and explanation generation, addressing the limitations of prior works.
  • We introduce the first application of Group Relative Policy Optimization (GRPO) [12] in this domain, enhancing explanation quality in multimodal, bilingual scenarios.
  • Our framework achieves new state-of-the-art performance on BiMiBench, substantially improving classification, localization, and explanation metrics over competitive baselines.
2 Related Work
2.1 Misinformation Detection
Datasets. Early misinformation detection datasets can be grouped into three main categories. Text-only datasets: FEVER [13] and FakeNewsNet [14] focus exclusively on textual fact verification and lack any visual modality, limiting their applicability to multimodal scenarios. Basic multimodal datasets: Fakeddit [15], MAIM [16], and EMU [17] introduce image-text pairs, but are typically limited to monolingual (English) content, coarse visual edits, and offer no explicit supervision for localization or explanation. Semantically aligned multimodal datasets: More recent benchmarks such as NewsCLIPpings [18], COSMOS [19], MMFakeBench [8] and DGM4 [9] move toward finer-grained multimodal reasoning. However, they remain focused on caption-style inputs and overlook subtitle-level or cross-lingual inconsistencies.
To address these limitations, we introduce BiMiBench, a large-scale benchmark curated from real-world news content that captures both visual and textual manipulations across five fine-grained misinformation types. As shown in Table 2.1, BiMiBench is the only benchmark that simultaneously supports bilingual subtitles, region-level localization, and both visual and textual tampering, offering the richest annotation scope among existing datasets. To evaluate the quality of visual manipulations, Figure 2 compares BiMiBench with recent benchmarks using a perceptual realism analysis across seven dimensions (e.g., content, composition, color), demonstrating that BiMiBench achieves significantly higher realism and diversity.
Table 1: Comparison of misinformation datasets. BiMiBench uniquely supports manipulations based on bilingual subtitles and localized content.
Dataset Multi- Sub- Bi- Textual Visual Local- News Mask
modal title lingual Manip. Manip. ization Domain
FEVER [13]
FakeNewsNet [14]
Fakeddit [15]
MAIM [16]
EMU [17]
NewsCLIPpings [18]
COSMOS [19]
DGM4 [9]
MMFakeBench [8]
BiMiBench (Ours)
Figure 2: Quality comparison of manipulated images across datasets, by the method of [20].
Methods. Current methods for multimodal misinformation detection typically focus on aligning image and text pairs, with an emphasis on out-of-context detection [21, 22, 9, 23, 24, 25]. NewsCLIPpings [18] adopts CLIP-based matching for image-caption congruence. COSMOS [19] employs self-supervised contrastive learning on out-of-context pairs but lacks supervision for fine-grained or localized manipulations. SNIFFER [7] and EMU [17] improve explainability via MLLMs and intent reasoning, yet rely on clean, monolingual captions. HAMMER [9] localizes manipulations through grounding or matching, but does not address subtitle-level or multilingual inconsistencies. MMD-Agent [8] broadens evaluation with mixed-source samples, though without explicit modeling of bilingual or region-specific manipulations. CroMe [26] introduces cross-modal metric learning over image, text, and fused representations, but remains limited to English captions. In contrast, we are the first to detect misinformation involving tampered images and bilingual subtitles, with support for localization, cross-lingual reasoning, and explanation.
2.2 Multi-modal Large Language Models
Model. Multimodal large language models (MLLMs) [27, 4, 28, 29, 30] integrate vision encoders with pretrained language models to enable unified visual-text reasoning. Flamingo [31] introduces prompt-based multimodal learning, showing strong few-shot results on tasks like VQA. BLIP-2 [28], MiniGPT-4 [29], and LLaVA series [30, 32, 33] further enhance visual instruction tuning, aligning modalities through instruction-following data to support tasks such as VQA, captioning, and dialogue. Recent models like InternVL3 [27], Qwen3-VL [34], and Gemini [35] improve generalization, multilinguality, and open-ended tool use. Despite these advances, most MLLMs focus on grounded understanding or generation and lack supervision for detecting fine-grained cross-modal inconsistencies—critical for misinformation detection [36, 37]. To improve efficiency, many lightweight MLLMs have emerged, such as Gemma [38]; in our work, we adopt Gemma 3 [39] as the base model to balance performance and scalability.
GRPO. Reinforcement learning (RL) has been shown to enhance reasoning in LLMs [40], with DeepSeek-R1 [5, 12] introducing Group Relative Policy Optimization (GRPO) to further improve explanation-driven multimodal performance. GRPO optimizes the relative ranking of outputs within grouped samples, enabling fine-grained discrimination between aligned and misaligned multimodal inputs. However, its potential remains underexplored in misinformation detection tasks involving subtle visual edits and bilingual subtitle inconsistencies.
Prior work [7, 8, 9] rarely tackles misinformation that involves both manipulated images and bilingual subtitles, and often lacks the ability to localize edits, reason across modalities, and adapt to new events. To address these gaps, we propose a unified framework that integrates bilingual visual-text reasoning, manipulation localization, explanation generation, and retrieval-based adaptation.
3 BiMiBench: A Benchmark for Bilingual Multimodal Misinformation
Existing benchmarks for multimodal misinformation primarily focus on monolingual captions or synthetic mismatches, lacking realism, cross-lingual scope, and fine-grained supervision. In real-world scenarios, misinformation often involves tampered images with bilingual (Chinese-English) subtitles, where inconsistencies may occur in any modality. To address this gap, we introduce BiMiBench, the first benchmark that supports not only bilingual multimodal reasoning but also manipulation, localization, and natural language explanation, enabling comprehensive evaluation in realistic settings.
3.1 Dataset Construction
Details. BiMiBench contains 104,000 samples derived from real-world news content. Specifically, we construct our dataset based on the VisualNews [41] dataset and select the portion sourced from the Washington Post to ensure factual grounding and linguistic consistency. Each sample includes an image paired with Chinese-English subtitles, with misinformation introduced through localized image editing, Chinese subtitle modification, or English subtitle modification. Approximately 80% of the samples contain manipulated content in at least one modality, while the remaining 20% are clean control examples. Image resolutions range from 640×480 to 1024×768 pixels, and subtitles are provided in both English and Chinese to support bilingual misinformation detection. BiMiBench captures cross-modal inconsistencies, offering a challenging testbed for evaluating multimodal and bilingual reasoning capabilities.
Construction. Our benchmark is constructed based on VisualNews, a large-scale image-text news dataset.
Figure 3: The data generation workflow used in constructing the BiMiBench benchmark.
To generate realistic misinformation samples, we design a multi-step data generation pipeline. Let $\mathcal{D}_{\text{orig}}=\{(I_{i},C_{i})\}_{i=1}^{N}$ denote the original dataset from VisualNews, where $I_{i}$ is an image and $C_{i}$ is its associated English caption.
Step 1: Textual Manipulation. Given an image-caption pair $(I_{i},C_{i})$ from VisualNews [41], we use the Gemma 3 [39] model $f_{\text{Gemma}}$ to generate a manipulated caption $\tilde{C}_{i}=f_{\text{Gemma}}(I_{i},C_{i},\text{PROMPT})$, introducing factual inconsistencies while preserving contextual plausibility.
Step 2. Bilingual Translation. We translate the original and manipulated English captions into Chinese using a translation API [42] $\mathcal{T}$, yielding bilingual subtitle pairs: $C^{\text{zh}}_{i}=\mathcal{T}(C_{i}),\tilde{C}^{\text{zh}}_{i}=\mathcal{T}(\tilde{C}_{i})$.
Step 3. Semantic Segmentation. To enable targeted visual editing, we use an instruction-based segmentation model LISA [43]:$f_{\text{LISA}}$ to extract object masks based on the manipulated caption: $M_{i}=f_{\text{LISA}}(I_{i},\tilde{C}_{i},\text{PROMPT})$, where $M_{i}$ highlights regions corresponding to entities in $\tilde{C}_{i}$.
Step 4. Visual Manipulation. We apply a visual manipulation function $\mathcal{V}$ to perform localized editing on the original image using the object masks: $\tilde{I}_{i}=\mathcal{V}(I_{i},M_{i})$, where the edited image $\tilde{I}_{i}$ semantically aligns with the manipulated caption $\tilde{C}_{i}$. To ensure visual diversity and realism, we adopt a mix of recent state-of-the-art image editing techniques, including FLUX [44], VAR [45], and SDXL [46].
Step 5. Label Assignment. For each final sample, we randomly select real or manipulated versions of the image and subtitles to construct a multimodal example: $S_{i}^{*}=S\left(I_{i}^{*},C_{i}^{\text{en}},L_{i}^{\text{zh}}\right),I_{i}^{*}\in\{I_{i},\tilde{I}_{i}\},C_{i}^{\text{en}}\in\{C_{i},\tilde{C}_{i}\},L_{i}^{\text{zh}}\in\{C^{\text{zh}}_{i},\tilde{C}^{\text{zh}}_{i}\}$, where misinformation may appear in any modality individually or in combination.
Table 2: BiMiBench Category Definitions.
Category Definition
All Consistent $I_{i}^{*}=I_{i},C_{i}^{\text{en}}=C_{i},L_{i}^{\text{zh}}=C_{i}^{\text{zh}}$
Image Manipulated $I_{i}^{*}=\tilde{I}_{i},C_{i}^{\text{en}}\neq\tilde{C}_{i}$ or $L_{i}^{\text{zh}}\neq\tilde{C}_{i}^{\text{zh}}$
Both Misaligned $I_{i}^{*}=I_{i},C_{i}^{\text{en}}=\tilde{C}_{i},L_{i}^{\text{zh}}=\tilde{C}_{i}^{\text{zh}}$
Chinese Misaligned $I_{i}^{*}=I_{i},C_{i}^{\text{en}}=C_{i},L_{i}^{\text{zh}}=\tilde{C}_{i}^{\text{zh}}$
English Misaligned $I_{i}^{*}=I_{i},C_{i}^{\text{en}}=\tilde{C}_{i},L_{i}^{\text{zh}}=C_{i}^{\text{zh}}$
All Inconsistent $I_{i}^{*}=\tilde{I}_{i},C_{i}^{\text{en}}=\tilde{C}_{i},L_{i}^{\text{zh}}=\tilde{C}_{i}^{\text{zh}}$
Each sample in BiMiBench is assigned to one of six mutually exclusive categories based on image manipulation and semantic alignment between the image $I_{i}^{*}$, English subtitle $C_{i}^{\text{en}}$, and Chinese subtitle $L_{i}^{\text{zh}}$. The categories cover typical multimodal misinformation patterns: (1) All Consistent: Image and both subtitles are consistent. (2) Image Manipulated: The image is tampered, and at least one subtitle does not align with it. (3) Both Misaligned: Image is real, but both subtitles are misleading. (4) Chinese Misaligned: Image is real, English subtitle is faithful, Chinese is misleading. (5) English Misaligned: Image is real, Chinese subtitle is faithful, English is misleading. (6) All Inconsistent: Image is tampered; both subtitles are misleading. Table 2 provides formal definitions for each class. All samples are manually reviewed for annotation quality and consistency. This design enables structured supervision of complex multimodal misinformation scenarios. Additional examples and detailed analysis of BiMiBench are provided in the Appendix.
4 BiMi: A Bilingual Multimodal Misinformation Detection Framework
Real-world multimodal misinformation often involves image edits and inconsistencies between bilingual subtitles. Existing models struggle to detect such cross-modal manipulations and lack adaptability to emerging events. We propose BiMi, a bilingual multimodal framework that localizes manipulated regions, detects cross-modal inconsistencies, and generates explanations. To enhance generalization, BiMi integrates an online retrieval module that provides real-time external context.
Figure 4: The overview of the training strategy. Three stages: domain alignment on news data, instruction tuning with task-specific prompts, and GRPO optimization with structured rewards. Each stage builds upon the previous one, progressively enhancing task performance.
4.1 Framework
Modeling. We adopt Gemma 3 [39] as the pre-trained model of BiMi due to its strong multilingual understanding and cross-modal reasoning capabilities. Gemma 3 is a multilingual MLLM with strong vision-language alignment and cross-lingual reasoning, making it well-suited for detecting subtle manipulations across images and bilingual subtitles. In our framework, the input image is overlaid with Chinese and English subtitles to form a unified visual input, which is then processed by a vision encoder to extract patch-level embeddings. These representations are fused through learned attention layers to form a unified multimodal context. This design allows the model to jointly reason over visual and bilingual textual cues, making it well-suited for detecting subtle manipulations such as object removal, semantic shifts, and cross-lingual inconsistencies.
Retrieval Module. To improve adaptability to emerging misinformation, BiMi integrates a retrieval module at inference time. We first extract Chinese and English subtitles from the image using OCR, then query the Google Search API [42] and collect the top-3 results. Retrieved texts are appended to the input prompt, providing up-to-date context that helps the model assess semantic consistency. This retrieval-augmented setup enables BiMi to generalize better to previously unseen or rapidly evolving misinformation cases.
4.2 Multi-stage Training
To effectively adapt the pretrained Gemma to the task of multilingual multimodal misinformation detection, we adopt a three-stage training strategy aimed at progressively aligning the model with the news domain, task-specific instructions, and high-quality reasoning objectives. Figure 4 shows the overview of the training strategy.
Stage 1: Domain Alignment. To adapt the model to the linguistic and visual characteristics of the news domain, we perform instruction-based tuning using image-caption pairs from VisualNews [41]. Each instance consists of an instruction prompt (e.g., Describe this image in English/Chinese) and a corresponding caption in either English or Chinese. We use captions from the Washington Post subset and translate each English caption into Chinese to ensure cross-lingual coverage. The model learns to generate captions conditioned on both the image and the prompt, thereby establishing strong visual-textual grounding across both languages. This stage enhances the model’s ability to handle real-world journalistic content and supports bilingual reasoning in downstream misinformation detection tasks.
Stage 2: Instruction Tuning. We fine-tune the model with single-turn instruction-following data tailored for multimodal misinformation detection. Each sample includes an image $X_{v}$, a bilingual subtitle pair, and a prompt $X_{\text{q}}$ that asks the model to perform multiple sub-tasks: detect manipulation, assess cross-modal consistency, and explain the decision. The assistant response $X_{\text{a}}$ includes structured answers and a natural language explanation. We follow a unified sequence format:
[…]
where $X_{\text{q}}$ contains the natural language instruction and bilingual subtitles, and $X_{\text{a}}$ consists of three binary labels and a free-form explanation, formatted as:
[…]
The model is trained to generate $X_{\text{a}}$ conditioned on $X_{\text{q}}$ and $X_{v}$, by maximizing the likelihood:
$P(X_{\text{a}}\mid X_{\text{q}},X_{v})=\prod_{i=1}^{L}P_{\theta}(x_{i}\mid x_{<i})$ (1)
where $x_{i}$ denotes the $i$-th token in the assistant response. Only the tokens in $X_{\text{a}}$ are used to compute the loss. This tuning step encourages the model to align with the structure and reasoning required for fine-grained misinformation detection.
Stage 3: GRPO-based Reasoning Optimization. To improve the model’s reasoning quality and explanation consistency, we apply Group Relative Policy Optimization (GRPO) [12], a reinforcement learning strategy that adjusts the policy by comparing outputs within a batch rather than relying on absolute rewards alone.
Given a question $q$, GRPO samples $N$ candidate responses $\{r_{1},r_{2},\ldots,r_{N}\}$ from the policy $\pi_{\theta}$ and evaluates each response $o_{i}$ using a reward function $R\left(q,o_{i}\right)$, which measures the quality of the candidate in the context of the given question. GRPO encourages the model to generate responses with higher advantages within the group by updating the policy $\pi_{\theta}$ using the following objective:
$\mathcal{J}_{\mathrm{GRPO}}(\theta)=\mathbb{E}\left[{\{o_{i}\}_{i=1}^{N}\sim\pi_{\theta_{\text{old}}}(q)}\right]\frac{1}{N}\sum_{i=1}^{N}\left\{\min\left[s_{1}\cdot A_{i},\;s_{2}\cdot A_{i}\right]-\beta\mathbb{D}_{\mathrm{KL}}\left(\pi_{\theta}\|\pi_{\mathrm{ref}}\right)\right\}$ (2)
$A_{i}$ $=\frac{r_{i}-\mathrm{mean}\{r_{1},r_{2},\ldots,r_{N}\}}{\mathrm{std}\{r_{1},r_{2},\ldots,r_{N}\}},\quad s_{1}=\frac{\pi_{\theta}(o_{i}\mid q)}{\pi_{\theta_{\text{old}}}(o_{i}\mid q)},\quad s_{2}=\mathrm{clip}\left(s_{1},\;1+\epsilon,\;1-\epsilon\right),$ (3)
where $A_{i}$ represents the advantage of the candidate response $o_{i}$ relative to other sampled responses. Following DeepSeek-R1, we use both format and accuracy reward.
Reward function. To optimize the reasoning ability and explanation quality of BiMi during the final training stage, we design a composite reward function tailored for GRPO. In our setting, we focus on three core tasks: misinformation classification, tampered region localization, and natural language explanation generation.
Format reward. To enforce structured outputs, we define a format reward $R_{\text{format}}$ that checks whether the model response follows a predefined format, using <answer></answer> tags for predictions and an optional <think></think> section for explanations. This ensures consistency across outputs and simplifies automatic evaluation. We use regular expressions for validation: $R_{\text{format}}=\begin{cases}0,&\text{if output matches expected format}\\ 1,&\text{otherwise}\end{cases}$.
Localization reward. We use an IoU-based reward $R_{\text{Loc}}$ to assess the alignment between predicted and ground-truth tampered regions. To simplify computation, both are represented as bounding boxes. The reward is defined as: $R_{\text{Loc}}=\frac{|\mathcal{M}_{\text{pred}}\cap\mathcal{M}_{\text{gt}}|}{|\mathcal{M}_{\text{pred}}\cup\mathcal{M}_{\text{gt}}|}$, where $\mathcal{M}_{\text{pred}}$ and $\mathcal{M}_{\text{gt}}$ denote the predicted and ground-truth bounding boxes, respectively.
Classification reward. To encourage accurate decision-making in multimodal misinformation detection, we define a classification reward $R_{\text{cls}}$ that compares the model’s predicted label with the ground-truth annotation. This applies to tasks such as detecting whether the image is manipulated or whether the subtitles are semantically consistent. $R_{\text{cls}}=\begin{cases}1,&\text{if }C_{\text{pred}}=C_{\text{gt}}\\ 0,&\text{otherwise}\end{cases}$.
The final reward combines all task-specific objectives, including formatting, classification, localization: $R_{\text{total}}=R_{\text{format}}+R_{\text{cls}}+R_{\text{loc}}$. This unified reward encourages the model to generate structured, accurate, and interpretable predictions across modalities.
This progressive training strategy equips BiMi with the capability to perform fine-grained classification, localization, and explanation, and is designed to support generalization to challenging, real-world misinformation cases.
5 Experiments
5.1 Experimental Setup
We evaluate BiMi on BiMiBench and MMFakeBench [8]. On BiMiBench, the model performs six-class classification, tampered region localization, and explanation generation. For MMFakeBench, which lacks subtitle information, we follow its four-class setting and provide text context via prompts.
Implementation Details. We use the Gemma 3 as the LLM, applying our three-stage post-training pipeline. All experiments are run on 48GB V100 GPUs with a batch size of 16, learning rate $1\times 10^{-5}$, and frozen vision encoder. GRPO is applied for two additional epochs. We design structured instruction prompts to guide the model in tasks, as detailed in Table 3.
Please analyze the given image containing both Chinese and English subtitles and complete the following three tasks: (1) Classification Task: classify the alignment between the image and the subtitles into one of the following six categories: "all consistent", "image manipulated", "both subtitles misaligned with image", "only English aligned", "only Chinese aligned", "all inconsistent". (2) Manipulation Detection: if the image has been manipulated, return one or more bounding boxes for the manipulated regions in the format: {"bbox":[x_min, y_min, x_max, y_max]}. If no manipulation is found, return an empty list. (3) Decision Explanation: briefly explain your thinking before the classification and any detected regions. Return your output using the following format, wrapped in tags:
<think>Your explanation here.</think>
<answer>{"classification": "result", "region": [{"bbox": [x_min, y_min, x_max, y_max]}]}</answer>
Table 3: Prompt templates used in BiMi for multimodal misinformation detection
Baselines. We compare BiMi against a range of strong multimodal baselines, including InternVL3 [27], Qwen3 [47], LLaVA-1.6 [33], and LLama3.1 [48], which represent leading approaches in vision-language modeling. We also include specialized misinformation detection systems such as SNIFFER [7] and MMD-Agent [8], designed for out-of-context and multimodal misinformation. All models are evaluated using the same inputs: the image with overlaid Chinese-English subtitles and a task-specific prompt. For ablation, we test BiMi variants without GRPO, without retrieval, and with alternative MLLM backbones.
Evaluation Metrics. We report Accuracy (ACC) and F1 score for classification, IoU for localization. Inspired by [30], we evaluate explanation quality via BERTScore [49] between response and reference explanations generated by prompting GPT-4o [50] with labels and manually verified for correctness.
Table 4: Main results on BiMiBench. BiMi achieves the best performance across all metrics. Bold and underline denote best and second-best results.
Method LLM Res. ACC F1 IoU BERT
InternVL3 [27] Qwen2.5-7B 224 20.85 18.42 7.23 71.82
InternVL3 [27] Qwen2.5-14B 224 28.47 26.71 14.38 76.54
Qwen3 [47] Qwen3-8B 224 21.43 14.96 6.28 73.92
LLaVA-1.6 [33] Vicuna-7B 224 22.91 20.74 8.91 75.77
LLama3.1 [48] Llama-3.1-8B 224 24.61 21.94 5.27 72.91
Gemma3 [48] Gemma3-4B-IT 336 17.48 11.73 4.29 70.84
SNIFFER [7] Vicuna-13B 224 37.90 33.63 80.41
MMD-Agent [8] Qwen2.5-7B 224 34.78 29.48
BiMi Gemma3-4B-IT 336 46.80 42.79 30.24 82.90
5.2 Main Results
Table 5: MMFB results.
Method ACC F1
InternVL3 42.69 34.80
Qwen3 44.82 39.49
LLaVA-1.6 30.09 24.31
LLama3.1 32.67 28.30
Gemma3 27.19 22.04
SNIFFER 67.49 61.44
MMD-Agent 62.78 52.83
BiMi 70.94 62.49
Performance Comparison. We compare BiMi with strong baselines across three subtasks on the BiMiBench test set: (1) misinformation detection, (2) tampered region localization, and (3) explanation generation. Results are summarized in Table 4. BiMi significantly outperforms all baselines in both all tasks. In particular, it achieves a +8.9 accuracy gain in image-subtitle consistency, a +15.9 localization accuracy, and a +2.5 BERTScore improvement in explanation generation over the strongest baseline. BiMi’s performance also benefits from the GRPO reward design, which improves explanation specificity and localization precision. The results on MMFakeBench (4-class classification task) are shown in Table 5.
Ablation Study. We investigate the contributions of each stage in our pipeline using an ablation study. (i) Domain Alignment. Removing this stage reduces accuracy by 2.35 and BERTScore by 4.48, indicating its role in grounding visual-textual reasoning on news content. (ii) Instruction Tuning. Eliminating instruction tuning leads to a drastic performance drop, accuracy falls by 31.18 and BERTScore by 12.06, highlighting its importance in enabling structured bilingual understanding. (iii) GRPO. Without GRPO, the model fails to generate specific explanations tied to image or
Table 6: Table: Ablation results (%). Red text shows a performance drop from the best variant.
Design Accuracy BERTScore
Full model 48.60 82.90
w/o Domain Align 46.25 (-2.35) 78.42 (-4.48)
w/o Instr. Tune 17.42 (-31.18) 70.84 (-12.06)
w/o GRPO 42.90 (-5.70) 74.81 (-8.09)
Answer First 46.66 (-1.94) 81.79 (-1.11)
subtitle content, reducing accuracy by 5.7 and BERTScore by 8.09. (iv) Output Order. Reversing the output order (classify before explain) slightly lowers accuracy (–1.94) and BERTScore (–1.11), suggesting that generating explanations first facilitates more coherent reasoning. Table 6 summarizes the results. These findings confirm that all components are essential, with instruction tuning and GRPO being particularly critical for boosting accuracy and explanation quality.
5.3 Explainability Analysis
Beyond accurate classification and localization, generating high-quality natural language explanations is crucial for interpretable misinformation detection. We analyze explanation ability through (1) qualitative comparisons, (2) the influence of the retrieval module, and (3) evaluation under bilingual.
Explanation Quality. To understand BiMi’s reasoning ability, we compare its output with InternVL3 [27] on a representative case (Figure 5). Both models recognize that the subtitles are semantically aligned, showing bilingual understanding. However, InternVL3 relies on external factual cues and overlooks the mismatch between image and text. In contrast, BiMi grounds its decision in the image, correctly identifies the manipulation, and localizes the tampered region. It also provides a clearer, step-by-step explanation, showing stronger visual grounding and more faithful reasoning.
Figure 5: Comparison of explanations from InternVL3 (middle) and our model (right). The top left is input. The bottom left is the original sample. Note: Due to space constraints, some model responses are truncated. Full examples and additional results are provided in the Appendix.
Effect of Retrieval. We analyze a real-world social media post with subtle bilingual inconsistency (Figure 6). The removed "student" in the Chinese translation alters the meaning. Without retrieval, the model misclassifies due to a lack of context. The retrieval result provides a key clue for classification. BiMi correctly identifies the inconsistency, showing its benefit for nuanced cross-lingual reasoning.
Table 7: Accuracy and BERTScore for each language variant (4-category classification).
Input Variant ACC BERTScore
CN-only 72.32
EN-only 82.48 83.44
Bilingual multimodal capability. To evaluate bilingual understanding, we construct a variant of the test set where the model receives only one subtitle, Chinese or English. This reduces the task to a 4-category classification. By ignoring cross-lingual consistency, the single-language setting reduces reasoning difficulty. As shown in Table 7, the model performs better with English-only inputs, likely due to the predominance of English data during training. This highlights a gap in bilingual support and the need for stronger Chinese-language modeling in this domain.
Figure 6: Social media example with subtle bilingual inconsistency. The Chinese translation omits "student", changing the meaning and potentially misleading Chinese readers.
Failure Cases. We observe that BiMi occasionally fails when manipulations are extremely subtle or when OCR fails to extract accurate subtitles. In such cases, either the localization is off-target, or the explanation is overly general. More failure cases and analysis are provided in the Appendix.
6 Conclusion and Limitation
Conclusion. We study the task of detecting multimodal inconsistencies in news images with bilingual (Chinese-English) subtitles. To support this, we introduce BiMiBench, a large-scale benchmark with 104K samples featuring realistic visual edits and cross-lingual mismatches. We propose BiMi, a bilingual multimodal framework combining a three-stage post-training strategy and retrieval-augmented reasoning. Experiments show that BiMi achieves state-of-the-art performance in classification, localization, and explanation, highlighting its effectiveness in real-world scenarios.
Limitation and future work. While BiMi achieves strong results on bilingual multimodal detection, several limitations remain. First, the current setup focuses on Chinese-English subtitles; extending to more languages would enhance generalization. Second, localization is limited to bounding boxes, which may miss fine-grained manipulations—future work could explore pixel-level segmentation. Third, BiMiBench samples use fixed subtitle styles and editing templates; expanding visual diversity and subtitle formats would better reflect real-world cases. Finally, while BiMi offers solid performance, its reasoning and explanation ability can be further improved, especially in ambiguous scenarios.
Appendix
Section A provides details on the dataset, including: dataset statistics, misinformation generation strategy, example visualization, and manual verification.
Section B outlines implementation and experimental details, including: model architecture and training settings, hyperparameters.
Section C discusses the broader impact and ethical implications of our work.
Section D presents qualitative examples.
Section E analyzes the failure case.
Appendix A Dataset Details
Dataset Statistics. BiMiBench comprises 104,000 samples spanning six fine-grained misinformation categories. Figure 7 summarizes key dataset properties. The left panel shows the distribution of samples across categories, ensuring diverse coverage of manipulation types.
Figure 7: Dataset statistics of BiMiBench. Left: distribution across six misinformation categories. Center: English subtitle length distribution. Right: normalized area of manipulated regions (bounding boxes height).
Misinformation Generation Strategy. To simulate realistic multimodal misinformation, we generate manipulated subtitles by introducing subtle yet misleading semantic alterations. These edits preserve grammatical fluency and contextual plausibility while modifying key factual elements such as subjects, causal relations, locations, or numerical values. We further introduce bilingual inconsistencies by translating the manipulated English subtitles into Chinese, allowing semantic shifts to emerge naturally during translation. This strategy enables diverse and challenging misinformation scenarios that reflect real-world cross-modal and cross-lingual manipulations. We design task-specific prompts to guide misinformation generation, ensuring fluency and semantic plausibility while introducing factual inconsistencies; an example prompt used in this process is shown in Table 8.
<system> Based on the image, create a fictional, sensational news headline or summary in one sentence. The story must be clearly fake but sound plausible, and must be no more than 30 words. The news should sound outrageous, scandalous, or unbelievable, yet still written in a serious, news-like tone. </system>
Table 8: Prompt templates used in BiMi for multimodal misinformation detection
Example Visualization. Figure 8 illustrates representative BiMiBench samples, covering various manipulation types, language inconsistencies, and corresponding ground truth labels.
Figure 8: Sample BiMiBench examples with bilingual subtitles and visual or text manipulations.
Manual Verification. To ensure data quality and annotation reliability, all samples in BiMiBench were manually verified by a team of five trained annotators, each with backgrounds in journalism, linguistics, or media studies, and prior experience with fact-checking or misinformation analysis. Annotators received targeted instruction on identifying realistic image manipulations, assessing semantic consistency across English-Chinese subtitle pairs, and applying fine-grained manipulation categories. During the review process, they checked the accuracy of tampered regions, validated bounding box alignment, and evaluated bilingual subtitles for cross-lingual fidelity and plausibility. Samples with OCR errors, ambiguous edits, mistranslations, or low visual quality were discarded. This multi-layered, expert-driven verification process ensures that BiMiBench maintains high annotation quality, semantic precision, and supports reliable benchmarking for multilingual multimodal misinformation detection.
Appendix B Implementation Details
Model Architecture. Our framework is built upon the publicly released Gemma 3 multimodal language model. The vision encoder is initialized from a pretrained ViT-L model, and remains frozen throughout all training stages. LoRA parameters are fine-tuned. The model accepts the inputs of the image with English and Chinese subtitle, which serialized into a unified prompt format. The image is encoded through a unified visual-language backbone.
Tokenization and Input Format. Both English and Chinese subtitles are tokenized using the SentencePiece tokenizer aligned with the Gemma 3 backbone. The input is formatted into structured prompts that specify the prediction task, marking different modalities. Input length is capped at 512 tokens.
Hyperparameter Summary. We fine-tune BiMi using the Gemma 3 model (4B) with a batch size of 8, a learning rate of $1\times 10^{-5}$, and a maximum input length of 512 tokens. The model is optimized with AdamW, using $\beta_{1}=0.9$, $\beta_{2}=0.98$, and a weight decay of 0.01. To stabilize training, we use gradient accumulation (4 steps) and mixed precision (FP16). The GRPO reward coefficient $\lambda$s are all set to 1. These hyperparameters are kept consistent across all three training stages unless otherwise specified.
Prompt Template. We design a structured instruction format to guide the model in reasoning across modalities and generating faithful explanations. The prompt template used during instruction tuning and inference is shown in Table 9. This format ensures consistent output structure, supporting multimodal reasoning (via <think>), interpretable classification, and region-level manipulation localization.
<system> You are an expert in misinformation detection area. A conversation between User and Assistant. The user asks a question, and the Assistant solves it. The assistant first thinks about the reasoning process in the mind and then provides the user with the answer. The reasoning process and answer are enclosed within <think> </think> and <answer> </answer> tags, respectively, i.e., <think> reasoning process here </think><answer> answer here </answer>. Assign one of the following six categories based on their mutual alignment: All Consistent: The image aligns well with both Chinese and English captions. No signs of manipulation or misalignment. Image Manipulated: The image is manipulated. Both captions truthfully describe the manipulated image. Both Misaligned: Both Chinese and English captions are manipulated. Neither caption correctly describes the image. Chinese Misaligned: Only the Chinese caption is manipulated. The English caption aligns correctly with the image. English Misaligned: Only the English caption is manipulated. The Chinese caption aligns correctly with the image. All Inconsistent: The image, Chinese caption, and English caption are all manipulated and mutually inconsistent. </system> <user> Please analyze the given image containing both Chinese and English subtitles and complete the following three tasks: (1) Classification Task: classify the alignment between the image and the subtitles into one of the following six categories: "all consistent", "image manipulated", "both subtitles misaligned with image", "only English aligned", "only Chinese aligned", "all inconsistent". (2) Manipulation Detection: if the image has been manipulated, return one or more bounding boxes for the manipulated regions in the format: {"bbox":[x_min, y_min, x_max, y_max]}. If no manipulation is found, return an empty list. (3) Decision Explanation: briefly explain your thinking before the classification and any detected regions. Return your output using the following format, wrapped in tags: <think> Your explanation here. </think> <answer> "classification": "result", "region": [{"bbox": [x_min, y_min, x_max, y_max]}] </answer> </user>
Table 9: Prompt templates used in BiMi for multimodal misinformation detection
Appendix C Broader Impact and Ethics
BiMi is designed to enhance the interpretability and reliability of multimodal misinformation detection, particularly in the context of bilingual news media. By providing both classification outputs and natural language explanations, BiMi promotes transparency in AI decision-making and supports fact-checkers, journalists, and platform moderators in identifying subtle manipulations across images and subtitles. To ensure ethical standards, all data samples in BiMiBench have been manually reviewed, linguistic clarity, and annotation consistency. Low-quality, ambiguous content has been excluded. The dataset construction avoids personally identifiable information and sensitive content.
BiMi is detection-focused and does not produce or promote fake content, reducing the risk of misuse. Nonetheless, we recognize that any publicly released model or dataset could potentially be repurposed. To mitigate this, we include usage guidelines and encourage responsible deployment, especially in downstream applications that may involve high-stakes decisions. We believe this work contributes positively by advancing research on robust, interpretable AI tools for combating harmful information while maintaining strong ethical safeguards throughout the development and release process.
Appendix D Qualitative Examples
Figure 9 10 11 12 13 14 15 16 show the examples.
Figure 9: Case of All Consistency.
Figure 10: Case of Chinese Misaligned.
Figure 11: Case of Image Manipulation.
Figure 12: Case of English Misaligned.
Figure 13: Case of Both Misaligned.
Figure 14: Case of English Misaligned.
Figure 15: Case of All Inconsistent.
Figure 16: Case of Both Misaligned.
Appendix E Failure Cases
Despite the overall effectiveness of our model, we observed several notable failure cases that highlight its current limitations. One recurring issue involves semantic misinterpretation, particularly when the input contains rare or ambiguous phrases. For example, in a prompt requiring multi-step reasoning or contextual understanding (e.g., “Describe the implication of climate change on coastal microeconomies”), the model occasionally generates generic or overly broad responses that fail to capture the specific nuances of the query. Another type of failure arises in edge cases involving domain-specific knowledge.
Figure 17 shows a failure case where the model misclassified an image manipulation example as “All Inconsistent.” The English caption aligns with the image, but the Chinese one introduces a false espionage narrative. Key issues include: misclassification; incorrect caption judgment—only the Chinese caption was wrong; missing bounding box. This highlights weaknesses in manipulation detection, multilingual judgment, and grounding.
Figure 17: Failure case

### v3

@@
 Abstract
-The increasing realism of multimodal content has made misinformation more subtle and more complex to detect, especially in news media, where images are frequently paired with bilingual subtitles (e.g., Chinese-English).
+The increasing realism of multimodal content has made misinformation more subtle and harder to detect, especially in news media where images are frequently paired with bilingual (e.g., Chinese-English) subtitles.
 Such content often includes localized image edits and cross-lingual inconsistencies that jointly distort meaning while remaining superficially plausible.
@@
 We further release BiMiBench, a large-scale and comprehensive benchmark constructed by systematically editing real news images and subtitles, comprising 104,000 samples with realistic manipulations across visual and linguistic modalities.
-To enhance interpretability, we apply Group Relative Policy Optimization (GRPO) to improve explanation quality, marking the first use of GRPO in this domain.
-Extensive experiments demonstrate that BiMi outperforms strong baselines by up to +8.9 in classification accuracy, +15.9 in localization accuracy, and +2.5 in explanation BERTScore, advancing state-of-the-art performance in realistic, multilingual misinformation detection.
-Code, models, and datasets are avaliable at https://github.com/RaffeLegend/misinformation.
+To enhance interpretability and cross-modal reasoning consistency, we apply Group Relative Policy Optimization (GRPO) to improve explanation quality, marking the first use of GRPO in this domain.
+Extensive experiments demonstrate that BiMi outperforms strong baselines by up to +8.9 in classification accuracy, +15.9 in localization accuracy, and +2.5 in explanation BERTScore, advancing state-of-the-art performance in realistic, bilingual misinformation detection.
+Code, models, and datasets will be released.
 1 Introduction
-The rapid advancement of generative models [1, 2, 3, 4, 5] has drastically lowered the barrier to producing highly realistic multimodal content [6].
-While enabling creative applications, this progress has also escalated the proliferation of misinformation that intricately blends visual and textual elements [7, 8].
-A particularly insidious form emerges in news media, where images are frequently overlaid with bilingual subtitles (e.g., Chinese-English).
-Here, subtle visual tampering and cross-lingual inconsistencies are exploited to distort facts while maintaining surface-level plausibility.
-This becomes especially dangerous for monolingual audiences, who may only comprehend one language and thus remain oblivious to manipulations in the other.
-A stark example occurred during early 2020, when a CDC report confirming the first U.S.
-case of COVID-19 community transmission [^1] https://edition.cnn.com/2020/02/27/health/us-cases-coronavirus-community-transmission/index.html was mistranslated by some Chinese social media as asserting that the virus originated in the United States.
-This misleading bilingual manipulation fueled significant public misunderstanding and geopolitical tensions.
-Such cases illustrate how multimodal misinformation increasingly leverages localized image edits and asymmetric subtitle translations to manipulate public perception across language communities.
-Addressing such sophisticated attacks demands more than conventional binary misinformation detection.
-It requires systems capable of region-level localization of manipulations, identification of cross-modal and cross-lingual inconsistencies, and generation of coherent, cross-lingual natural language explanations—capabilities that remain underexplored.
-Despite encouraging progress [7, 8, 9, 10, 11], existing systems exhibit three critical limitations.
-First, most rely on global image-text matching, lacking the fine-grained region-level localization necessary to expose subtle manipulations.
-Second, their explanations are often superficial and generic, providing little actionable insight into why content is misleading.
-Third, they neglect the modeling of bilingual subtitle inconsistencies, overlooking the semantic divergence arising from intentionally misleading translations.
-These shortcomings underline the need for a unified framework that reasons at a fine granularity across visual and bilingual textual modalities.
 Figure 1: Comparison of traditional vs.
@@
 Traditional tasks focus on visual-text consistency with limited outputs (left).
-Our setting uses tampered news images and bilingual subtitles, enabling richer outputs including region localization, cross-modal consistency, and explanation (right).
+Our setting uses tampered images and bilingual subtitles, enabling richer outputs including region localization, cross-modal consistency, and explanation (right).
 Red indicates error, green indicates correctness.
 Best viewed in color.
-However, building such a framework is technically challenging.
-Region localization demands precise visual-textual alignment that is beyond the capacity of most pretrained MLLMs.
-Generating faithful explanations in bilingual, multimodal settings requires accurate identification and articulation of inconsistencies across modalities and languages.
-Moreover, cross-lingual reasoning introduces unique complexities, as even fluent translations may introduce subtle but impactful semantic shifts that obscure the misinformation signal.
-These challenges are further exacerbated in the news domain, characterized by rapidly evolving, diverse, and context-dependent content.
-To tackle these challenges, we introduce BiMi, the first bilingual multimodal misinformation detection framework that seamlessly integrates region-level manipulation localization, cross-modal and cross-lingual consistency detection, and natural language explanation generation.
-An overview of our task and outputs is illustrated in Figure 1.
-To enhance generalization to emerging events, BiMi is equipped with an online retrieval module that augments model reasoning with real-time external knowledge.
-Furthermore, to facilitate systematic evaluation, we construct BiMiBench, the first large-scale benchmark for this setting, comprising 104,000 samples with realistic manipulations of news images and bilingual subtitles.
-Through extensive experiments, BiMi achieves state-of-the-art performance, surpassing strong baselines by +8.9 in classification accuracy, +15.9 in localization accuracy, and +2.5 in explanation BERTScore.
-Our contributions are summarized as follows:
-  • We release BiMiBench, the first large-scale benchmark for bilingual multimodal misinformation detection, containing 104K samples with fine-grained image manipulations and bilingual subtitle inconsistencies.
-  • We propose BiMi, a unified framework that detects image-subtitle misinformation via region localization, cross-modal consistency detection, and explanation generation, addressing the limitations of prior works.
-  • We introduce the first application of Group Relative Policy Optimization (GRPO) [12] in this domain, enhancing explanation quality in multimodal, bilingual scenarios.
-  • Our framework achieves new state-of-the-art performance on BiMiBench, substantially improving classification, localization, and explanation metrics over competitive baselines.
+The rapid progress of large generative models [achiam2023gpt, rombach2022high, zhao2025surveylargelanguagemodels, qwen, guo2025deepseek, hurst2024gpt] has substantially lowered the barrier to producing highly realistic multimodal content [openai2023dalle3], enabling new creative applications but also heightening the risk of multimodal misinformation, where manipulated images and accompanying text jointly mislead audiences [qi2024sniffer, liu2024mmfakebench].
+A recent UNESCO report warns that the widespread adoption of generative-AI tools can be exploited to fabricate convincing visual–textual narratives, thereby creating fertile ground for large-scale multimodal misinformation.
+Lost in Translation [quelle2025lost, guo2019future] further shows that misinformation frequently crosses language boundaries, underscoring the global risk of cross-lingual diffusion.
+A striking example occurred in early 2020, when a CDC report confirming the first U.S.
+case of COVID-19 community transmission was mistranslated on Chinese social media to claim that the virus originated in the United States, fueling public misunderstanding and geopolitical tension.
+These observations reveal how multimodal misinformation can exploit localized image edits and asymmetric subtitle translations to manipulate public perception across language communities [abdali2024multi].
+This creates an urgent need for methods capable of detecting fine-grained multimodal and cross-lingual inconsistencies—capabilities that remain largely underexplored.
+We therefore formulate the problem as bilingual multimodal misinformation detection: jointly localizing manipulated image regions and identifying cross-lingual inconsistencies while generating faithful natural-language explanations.
+Despite notable progress in multimodal misinformation detection [wu2019misinformation, tahmasebi2024multimodal, wan2024dell] and deepfake detection [huang2025so, huang2025sida, zhao2025deepfakebench, shao2025deepfake], existing approaches remain constrained by a predominant focus on coarse-grained image–text alignment and monolingual settings.
+First, many systems match images and captions only at the global level [qi2024sniffer, liu2024mmfakebench], leaving them unable to localize fine-grained, region-level manipulations such as localized edits or subtle subtitle tampering.
+Second, their explanation modules typically produce high-level, generic rationales [shao2025deepfake], providing little concrete evidence of why specific content is misleading.
+Third, no prior work has specifically addressed bilingual subtitle inconsistencies, leaving the semantic divergence introduced by deliberately misleading cross-lingual translations largely unexplored.
+These limitations collectively underscore the need for a unified framework that can jointly reason over visual content and bilingual text with precise localization and faithful explanation—capabilities that motivate the approach we develop in this work.
+However, building such a framework is technically demanding.
+Achieving fine-grained grounding between image regions and textual cues remains beyond the capability of most existing multimodal large language models (MLLMs).
+Moreover, producing faithful explanations in a bilingual multimodal setting requires accurate detection and articulation of inconsistencies across both modalities and languages.
+Cross-lingual reasoning is further complicated by subtle semantic shifts that even fluent translations can introduce, often obscuring misinformation signals.
+These challenges are especially acute in the news domain, where content evolves rapidly and exhibits high diversity and context dependence, making it non-trivial to design a system that jointly delivers precise localization, robust cross-lingual reasoning, and faithful explanation.
+To address these challenges, we introduce BiMi, the first framework that jointly targets bilingual subtitles inconsistency, region-level manipulation localization, and natural-language explanation generation.
+Figure 1 compares our unified task setup with prior multimodal approaches.
+To improve generalization to emerging events, BiMi incorporates an online retrieval module that augments model reasoning with real-time external knowledge.
+We further construct BiMiBench, the first large-scale benchmark for this setting, containing 104,000 news-image samples with realistic manipulations of visual content and bilingual subtitles.
+Extensive experiments demonstrate that BiMi establishes a new state of the art, outperforming strong baselines by +8.9 in classification accuracy, +15.9 in localization accuracy, and +2.5 in explanation BERTScore, demonstrating robust overall improvements.
+Our main contributions are as follows:
+  • BiMiBench: the first large-scale benchmark for bilingual multimodal misinformation detection, comprising 104K news–image samples with fine-grained visual manipulations and bilingual subtitle inconsistencies.
+  • BiMi Framework: a unified model that detects image–subtitle misinformation through region-level manipulation localization, multimodal and cross-lingual consistency detection, and natural-language explanation generation, directly addressing key limitations of prior work.
+  • GRPO for Explanation: the first application of Group Relative Policy Optimization (GRPO) [shao2024deepseekmath] to improve the quality and faithfulness of bilingual multimodal explanations.
+  • State-of-the-Art Results: BiMi achieves state-of-the-art performance on BiMiBench, with significant improvements in classification, localization, and explanation BERTScore over strong baselines.
 2 Related Work
@@
 Early misinformation detection datasets can be grouped into three main categories.
-Text-only datasets: FEVER [13] and FakeNewsNet [14] focus exclusively on textual fact verification and lack any visual modality, limiting their applicability to multimodal scenarios.
-Basic multimodal datasets: Fakeddit [15], MAIM [16], and EMU [17] introduce image-text pairs, but are typically limited to monolingual (English) content, coarse visual edits, and offer no explicit supervision for localization or explanation.
-Semantically aligned multimodal datasets: More recent benchmarks such as NewsCLIPpings [18], COSMOS [19], MMFakeBench [8] and DGM4 [9] move toward finer-grained multimodal reasoning.
-However, they remain focused on caption-style inputs and overlook subtitle-level or cross-lingual inconsistencies.
-To address these limitations, we introduce BiMiBench, a large-scale benchmark curated from real-world news content that captures both visual and textual manipulations across five fine-grained misinformation types.
-As shown in Table 2.1, BiMiBench is the only benchmark that simultaneously supports bilingual subtitles, region-level localization, and both visual and textual tampering, offering the richest annotation scope among existing datasets.
-To evaluate the quality of visual manipulations, Figure 2 compares BiMiBench with recent benchmarks using a perceptual realism analysis across seven dimensions (e.g., content, composition, color), demonstrating that BiMiBench achieves significantly higher realism and diversity.
-Table 1: Comparison of misinformation datasets. BiMiBench uniquely supports manipulations based on bilingual subtitles and localized content.
-Dataset Multi- Sub- Bi- Textual Visual Local- News Mask
-modal title lingual Manip. Manip. ization Domain
-FEVER [13]
-FakeNewsNet [14]
-Fakeddit [15]
-MAIM [16]
-EMU [17]
-NewsCLIPpings [18]
-COSMOS [19]
-DGM4 [9]
-MMFakeBench [8]
+Text-only: FEVER [Thorne18Fever] and FakeNewsNet [shu2020fakenewsnet] target textual fact verification and lack visual modality.
+Basic multimodal: Fakeddit [nakamura2019r], MAIM [jaiswal2017multimedia], and EMU [da-etal-2021-edited] pair images with text but are monolingual, use coarse edits, and provide no supervision for localization or explanation.
+Fine-grained multimodal: NewsCLIPpings [luo-etal-2021-newsclippings], COSMOS [aneja2021cosmos], MMFakeBench [liu2024mmfakebench], and DGM4 [shao2023detecting] enable richer reasoning yet still focus on caption-style inputs and ignore subtitle-level or cross-lingual inconsistencies.
+We introduce BiMiBench, which covers five types of visual and textual misinformation, uniquely supporting bilingual subtitles, region-level localization, and dual-modality tampering (Table 1); a perceptual analysis (Figure 2) shows it achieves higher realism and diversity than existing datasets.
+Table 1: Comparison of misinformation datasets. BiMiBench uniquely supports bilingual and localized manipulations.
+Dataset Sub. Bi. Text Vis. Loc. News Mask
+FEVER [Thorne18Fever]
+FakeNewsNet [shu2020fakenewsnet]
+Fakeddit [nakamura2019r]
+MAIM [jaiswal2017multimedia]
+EMU [da-etal-2021-edited]
+NewsCLIPpings [luo-etal-2021-newsclippings]
+COSMOS [aneja2021cosmos]
+DGM4 [shao2023detecting]
+MMFakeBench [liu2024mmfakebench]
 BiMiBench (Ours)
-Figure 2: Quality comparison of manipulated images across datasets, by the method of [20].
 Methods.
-Current methods for multimodal misinformation detection typically focus on aligning image and text pairs, with an emphasis on out-of-context detection [21, 22, 9, 23, 24, 25].
-NewsCLIPpings [18] adopts CLIP-based matching for image-caption congruence.
-COSMOS [19] employs self-supervised contrastive learning on out-of-context pairs but lacks supervision for fine-grained or localized manipulations.
-SNIFFER [7] and EMU [17] improve explainability via MLLMs and intent reasoning, yet rely on clean, monolingual captions.
-HAMMER [9] localizes manipulations through grounding or matching, but does not address subtitle-level or multilingual inconsistencies.
-MMD-Agent [8] broadens evaluation with mixed-source samples, though without explicit modeling of bilingual or region-specific manipulations.
-CroMe [26] introduces cross-modal metric learning over image, text, and fused representations, but remains limited to English captions.
-In contrast, we are the first to detect misinformation involving tampered images and bilingual subtitles, with support for localization, cross-lingual reasoning, and explanation.
+Current multimodal misinformation methods mainly align image–text pairs for out-of-context detection [przybyla2020capturing, qi2021improving, shao2023detecting].
+NewsCLIPpings [luo-etal-2021-newsclippings] and COSMOS [aneja2021cosmos] use CLIP-based or contrastive learning but lack supervision for fine-grained localization.
+SNIFFER [qi2024sniffer] and EMU [da-etal-2021-edited] add explainability via MLLMs yet rely on clean monolingual captions.
+HAMMER [shao2023detecting] localizes manipulations but ignores subtitle-level or multilingual cues.
+MMD-Agent [liu2024mmfakebench] and CroMe [choi2025cromemultimodalfakenews] broaden evaluation but remain limited to English captions.
+In contrast, our framework is the first to handle tampered images with bilingual subtitles, supporting localization, reasoning and explanation.
 2.2 Multi-modal Large Language Models
 Model.
-Multimodal large language models (MLLMs) [27, 4, 28, 29, 30] integrate vision encoders with pretrained language models to enable unified visual-text reasoning.
-Flamingo [31] introduces prompt-based multimodal learning, showing strong few-shot results on tasks like VQA.
-BLIP-2 [28], MiniGPT-4 [29], and LLaVA series [30, 32, 33] further enhance visual instruction tuning, aligning modalities through instruction-following data to support tasks such as VQA, captioning, and dialogue.
-Recent models like InternVL3 [27], Qwen3-VL [34], and Gemini [35] improve generalization, multilinguality, and open-ended tool use.
-Despite these advances, most MLLMs focus on grounded understanding or generation and lack supervision for detecting fine-grained cross-modal inconsistencies—critical for misinformation detection [36, 37].
-To improve efficiency, many lightweight MLLMs have emerged, such as Gemma [38]; in our work, we adopt Gemma 3 [39] as the base model to balance performance and scalability.
-GRPO.
-Reinforcement learning (RL) has been shown to enhance reasoning in LLMs [40], with DeepSeek-R1 [5, 12] introducing Group Relative Policy Optimization (GRPO) to further improve explanation-driven multimodal performance.
-GRPO optimizes the relative ranking of outputs within grouped samples, enabling fine-grained discrimination between aligned and misaligned multimodal inputs.
-However, its potential remains underexplored in misinformation detection tasks involving subtle visual edits and bilingual subtitle inconsistencies.
-Prior work [7, 8, 9] rarely tackles misinformation that involves both manipulated images and bilingual subtitles, and often lacks the ability to localize edits, reason across modalities, and adapt to new events.
-To address these gaps, we propose a unified framework that integrates bilingual visual-text reasoning, manipulation localization, explanation generation, and retrieval-based adaptation.
+Multimodal large language models (MLLMs) have greatly advanced cross-modal reasoning.
+Representative models such as Flamingo [alayrac2022flamingo], BLIP-2 [li2023blip], MiniGPT-4 [zhu2023minigpt4enhancingvisionlanguageunderstanding], Gemini [gemini2023, comanici2025gemini], InternVL-3 [zhu2025internvl3exploringadvancedtraining], Qwen-VL [bai2023qwenvlversatilevisionlanguagemodel, xu2025qwen2, qwen3], Gpt-4o [hurst2024gpt] and LLaVA [liu2023llava, liu2024llavanext] extend pretrained LLMs with cross-modal attention and instruction tuning for vision-language tasks.
+DeepSeek-R1 [guo2025deepseek, shao2024deepseekmath] further employs GRPO to enhance explanation quality.
+Yet most MLLMs focus on grounded understanding or generation [wu2025grounded, lin2025boosting] and are not trained to detect cross-modal inconsistencies.
+We adapt Gemma 3 [team2025gemma] to strengthen reasoning on visual and bilingual cues for misinformation detection.
+Training.
+MLLMs training typically includes large-scale pre-training followed by post-training with supervised fine-tuning (SFT) and reinforcement learning from human feedback (RLHF)[ouyang2022training].
+RLHF aligns outputs with human preferences using a reward model; methods such as Proximal Policy Optimization (PPO) [schulman2017proximal], Direct Preference Optimization (DPO) [rafailov2023direct], and GRPO [guo2025deepseek, shao2024deepseekmath] refine policies through preference ranking.
+RLHF remains explored for detecting visual edits and bilingual subtitle inconsistencies.
+Figure 2: Image quality comparison across datasets using the perceptual evaluation method of [chen2024evaluating].
+BiMiBench achieves higher visual quality across dimensions than prior datasets, indicating closer alignment with real-world social media imagery.
 3 BiMiBench: A Benchmark for Bilingual Multimodal Misinformation
@@
 In real-world scenarios, misinformation often involves tampered images with bilingual (Chinese-English) subtitles, where inconsistencies may occur in any modality.
-To address this gap, we introduce BiMiBench, the first benchmark that supports not only bilingual multimodal reasoning but also manipulation, localization, and natural language explanation, enabling comprehensive evaluation in realistic settings.
-3.1 Dataset Construction
+We introduce BiMiBench, a benchmark for multimodal misinformation detection with bilingual subtitles, providing labels and explanations for joint evaluation of classification, localization, and natural language explanation quality.
+3.1 Benchmark Construction
+Figure 3: The data generation workflow used in constructing the BiMiBench benchmark.
 Details.
-BiMiBench contains 104,000 samples derived from real-world news content.
-Specifically, we construct our dataset based on the VisualNews [41] dataset and select the portion sourced from the Washington Post to ensure factual grounding and linguistic consistency.
-Each sample includes an image paired with Chinese-English subtitles, with misinformation introduced through localized image editing, Chinese subtitle modification, or English subtitle modification.
-Approximately 80% of the samples contain manipulated content in at least one modality, while the remaining 20% are clean control examples.
-Image resolutions range from 640×480 to 1024×768 pixels, and subtitles are provided in both English and Chinese to support bilingual misinformation detection.
-BiMiBench captures cross-modal inconsistencies, offering a challenging testbed for evaluating multimodal and bilingual reasoning capabilities.
+BiMiBench comprises 104,000 real-world news samples derived from the VisualNews [liu-etal-2021-visual] corpus, a professionally curated collection of image–text pairs with broad topical diversity and reliable editorial quality.
+Each sample pairs an image with Chinese–English subtitles; misinformation is introduced via localized image edits or Chinese/English subtitle modifications.
+About 80% of the samples contain manipulations in at least one modality and 20% serve as clean controls.
+Images (640×480–1024×768) and bilingual subtitles provide a challenging testbed for evaluating multimodal and bilingual reasoning.
 Construction. Our benchmark is constructed based on VisualNews, a large-scale image-text news dataset.
-Figure 3: The data generation workflow used in constructing the BiMiBench benchmark.
-To generate realistic misinformation samples, we design a multi-step data generation pipeline.
-Let $\mathcal{D}_{\text{orig}}=\{(I_{i},C_{i})\}_{i=1}^{N}$ denote the original dataset from VisualNews, where $I_{i}$ is an image and $C_{i}$ is its associated English caption.
+To generate realistic misinformation samples, we design a multi-step data generation pipeline shown in Figure 3.
+Let $\mathcal{D}_{\text{orig}}=\{(I_{i},C_{i})\}_{i=1}^{N}$ denote the original dataset, where $I_{i}$ is an image and $C_{i}$ is its associated English caption.
 Step 1: Textual Manipulation.
-Given an image-caption pair $(I_{i},C_{i})$ from VisualNews [41], we use the Gemma 3 [39] model $f_{\text{Gemma}}$ to generate a manipulated caption $\tilde{C}_{i}=f_{\text{Gemma}}(I_{i},C_{i},\text{PROMPT})$, introducing factual inconsistencies while preserving contextual plausibility.
+Given an image-caption pair $(I_{i},C_{i})$ from VisualNews [liu-etal-2021-visual], we use the Gemma 3 [team2025gemma] model $f_{\text{Gemma}}$ to generate a manipulated caption $\tilde{C}_{i}=f_{\text{Gemma}}(I_{i},C_{i},\text{PROMPT})$, introducing inconsistencies while preserving contextual plausibility.
 Step 2.
 Bilingual Translation.
-We translate the original and manipulated English captions into Chinese using a translation API [42] $\mathcal{T}$, yielding bilingual subtitle pairs: $C^{\text{zh}}_{i}=\mathcal{T}(C_{i}),\tilde{C}^{\text{zh}}_{i}=\mathcal{T}(\tilde{C}_{i})$.
-Step 3.
-Semantic Segmentation.
-To enable targeted visual editing, we use an instruction-based segmentation model LISA [43]:$f_{\text{LISA}}$ to extract object masks based on the manipulated caption: $M_{i}=f_{\text{LISA}}(I_{i},\tilde{C}_{i},\text{PROMPT})$, where $M_{i}$ highlights regions corresponding to entities in $\tilde{C}_{i}$.
+We translate the original and manipulated English captions into Chinese using a translation API [googleapi2024] $\mathcal{T}$, selected for its accuracy and reliability in large-scale bilingual news translation, yielding bilingual subtitle pairs: $C^{\text{zh}}_{i}=\mathcal{T}(C_{i}),\tilde{C}^{\text{zh}}_{i}=\mathcal{T}(\tilde{C}_{i})$.
+[…]
 Step 4.
@@
 We apply a visual manipulation function $\mathcal{V}$ to perform localized editing on the original image using the object masks: $\tilde{I}_{i}=\mathcal{V}(I_{i},M_{i})$, where the edited image $\tilde{I}_{i}$ semantically aligns with the manipulated caption $\tilde{C}_{i}$.
-To ensure visual diversity and realism, we adopt a mix of recent state-of-the-art image editing techniques, including FLUX [44], VAR [45], and SDXL [46].
+To ensure visual diversity and realism, we adopt a mix of recent state-of-the-art image editing techniques, including FLUX [flux2024], VAR [tian2024visual], and SDXL [podell2023sdxlimprovinglatentdiffusion].
 Step 5.
@@
 Category Definition
-All Consistent $I_{i}^{*}=I_{i},C_{i}^{\text{en}}=C_{i},L_{i}^{\text{zh}}=C_{i}^{\text{zh}}$
-Image Manipulated $I_{i}^{*}=\tilde{I}_{i},C_{i}^{\text{en}}\neq\tilde{C}_{i}$ or $L_{i}^{\text{zh}}\neq\tilde{C}_{i}^{\text{zh}}$
-Both Misaligned $I_{i}^{*}=I_{i},C_{i}^{\text{en}}=\tilde{C}_{i},L_{i}^{\text{zh}}=\tilde{C}_{i}^{\text{zh}}$
-Chinese Misaligned $I_{i}^{*}=I_{i},C_{i}^{\text{en}}=C_{i},L_{i}^{\text{zh}}=\tilde{C}_{i}^{\text{zh}}$
-English Misaligned $I_{i}^{*}=I_{i},C_{i}^{\text{en}}=\tilde{C}_{i},L_{i}^{\text{zh}}=C_{i}^{\text{zh}}$
-All Inconsistent $I_{i}^{*}=\tilde{I}_{i},C_{i}^{\text{en}}=\tilde{C}_{i},L_{i}^{\text{zh}}=\tilde{C}_{i}^{\text{zh}}$
-Each sample in BiMiBench is assigned to one of six mutually exclusive categories based on image manipulation and semantic alignment between the image $I_{i}^{*}$, English subtitle $C_{i}^{\text{en}}$, and Chinese subtitle $L_{i}^{\text{zh}}$.
-The categories cover typical multimodal misinformation patterns: (1) All Consistent: Image and both subtitles are consistent.
-(2) Image Manipulated: The image is tampered, and at least one subtitle does not align with it.
-(3) Both Misaligned: Image is real, but both subtitles are misleading.
-(4) Chinese Misaligned: Image is real, English subtitle is faithful, Chinese is misleading.
-(5) English Misaligned: Image is real, Chinese subtitle is faithful, English is misleading.
-(6) All Inconsistent: Image is tampered; both subtitles are misleading.
-Table 2 provides formal definitions for each class.
-All samples are manually reviewed for annotation quality and consistency.
-This design enables structured supervision of complex multimodal misinformation scenarios.
-Additional examples and detailed analysis of BiMiBench are provided in the Appendix.
+All Consistent $I_{i}^{*}=I_{i},\,C_{i}^{\mathrm{en}}=C_{i},\,L_{i}^{\mathrm{zh}}=C_{i}^{\mathrm{zh}}$
+Image Manipulated $I_{i}^{*}=\tilde{I}_{i},\,C_{i}^{\mathrm{en}}\neq\tilde{C}_{i}$ or $L_{i}^{\mathrm{zh}}\neq\tilde{C}_{i}^{\mathrm{zh}}$
+Both Misaligned $I_{i}^{*}=I_{i},\,C_{i}^{\mathrm{en}}=\tilde{C}_{i},\,L_{i}^{\mathrm{zh}}=\tilde{C}_{i}^{\mathrm{zh}}$
+Chinese Misaligned $I_{i}^{*}=I_{i},\,C_{i}^{\mathrm{en}}=C_{i},\,L_{i}^{\mathrm{zh}}=\tilde{C}_{i}^{\mathrm{zh}}$
+English Misaligned $I_{i}^{*}=I_{i},\,C_{i}^{\mathrm{en}}=\tilde{C}_{i},\,L_{i}^{\mathrm{zh}}=C_{i}^{\mathrm{zh}}$
+All Inconsistent $I_{i}^{*}=\tilde{I}_{i},\,C_{i}^{\mathrm{en}}=\tilde{C}_{i},\,L_{i}^{\mathrm{zh}}=\tilde{C}_{i}^{\mathrm{zh}}$
+Categories.
+Each BiMiBench sample is labeled into one of six categories according to consistency among image $I_{i}^{*}$, English subtitle $C_{i}^{\text{en}}$, and Chinese subtitle $L_{i}^{\text{zh}}$.
+(1) All Consistent: all modalities agree; (2) Image Manipulated: image tampered, at least one subtitle mismatched; (3) Both Misaligned: image real, both subtitles misleading; (4) Chinese Misaligned: only Chinese subtitle misleading; (5) English Misaligned: only English subtitle misleading; (6) All Inconsistent: image tampered and both subtitles misleading.
+Formal definitions show in Table 2.
+All samples were manually reviewed to ensure annotation quality and consistency.
+Each item was independently checked by two annotators with a senior reviewer resolving disagreements, following predefined guidelines on factual correctness and translation fidelity.
+To address potential ethical and privacy concerns, we used only publicly available news content, removed any personally identifiable information, and release the dataset solely for non-commercial research in accordance with the source licenses.
+This design enables structured supervision of complex multimodal misinformation scenarios while maintaining high standards of data integrity and ethical compliance.
+Detailed information about the BiMiBench can be found in Appendix.
 4 BiMi: A Bilingual Multimodal Misinformation Detection Framework
@@
 To enhance generalization, BiMi integrates an online retrieval module that provides real-time external context.
-Figure 4: The overview of the training strategy.
-Three stages: domain alignment on news data, instruction tuning with task-specific prompts, and GRPO optimization with structured rewards.
-Each stage builds upon the previous one, progressively enhancing task performance.
+Figure 4: The overview of the training strategy. Three stages: domain alignment on news data, instruction tuning with task-specific prompts, and GRPO optimization with structured rewards.
 4.1 Framework
 Modeling.
-We adopt Gemma 3 [39] as the pre-trained model of BiMi due to its strong multilingual understanding and cross-modal reasoning capabilities.
-Gemma 3 is a multilingual MLLM with strong vision-language alignment and cross-lingual reasoning, making it well-suited for detecting subtle manipulations across images and bilingual subtitles.
-In our framework, the input image is overlaid with Chinese and English subtitles to form a unified visual input, which is then processed by a vision encoder to extract patch-level embeddings.
-These representations are fused through learned attention layers to form a unified multimodal context.
-This design allows the model to jointly reason over visual and bilingual textual cues, making it well-suited for detecting subtle manipulations such as object removal, semantic shifts, and cross-lingual inconsistencies.
+We adopt Gemma 3 [team2025gemma] as BiMi’s backbone for its strong multilingual understanding and vision–language alignment, enabling detection of subtle manipulations across images and bilingual subtitles.
+The input image with overlaid Chinese and English subtitles is encoded into patch-level embeddings, which are fused through attention layers to create a unified multimodal context.
+This design allows joint reasoning over visual and bilingual textual cues to identify object manipulation, semantic shifts, and cross-lingual inconsistencies.
 Retrieval Module.
-To improve adaptability to emerging misinformation, BiMi integrates a retrieval module at inference time.
-We first extract Chinese and English subtitles from the image using OCR, then query the Google Search API [42] and collect the top-3 results.
-Retrieved texts are appended to the input prompt, providing up-to-date context that helps the model assess semantic consistency.
-This retrieval-augmented setup enables BiMi to generalize better to previously unseen or rapidly evolving misinformation cases.
+To enhance adaptability to emerging misinformation, BiMi employs a retrieval module at inference.
+Chinese and English subtitles are extracted via OCR to form a bilingual query to the Google Search API [googleapi2024]; the top-3 retrieved passages are prepended to the input as a unified prompt $\mathcal{P}=\operatorname{concat}(R,I,S)$, where $R$, II, and $S$ denote the retrieved context, image, and subtitles.
+This auxiliary context supplies timely external information that improves the model’s generalization to previously unseen or rapidly evolving misinformation, while the final predictions remain grounded in the image–subtitle content.
 4.2 Multi-stage Training
@@
 Stage 1: Domain Alignment.
-To adapt the model to the linguistic and visual characteristics of the news domain, we perform instruction-based tuning using image-caption pairs from VisualNews [41].
-Each instance consists of an instruction prompt (e.g., Describe this image in English/Chinese) and a corresponding caption in either English or Chinese.
-We use captions from the Washington Post subset and translate each English caption into Chinese to ensure cross-lingual coverage.
-The model learns to generate captions conditioned on both the image and the prompt, thereby establishing strong visual-textual grounding across both languages.
-This stage enhances the model’s ability to handle real-world journalistic content and supports bilingual reasoning in downstream misinformation detection tasks.
+To adapt the model to the linguistic and visual traits of the news domain, we perform instruction-based tuning on VisualNews [liu-etal-2021-visual] image–caption pairs.
+Each instance uses an instruction prompt (e.g., Describe this image in English/Chinese) with an English or translated Chinese caption.
+This trains the model to generate captions from images and prompts, strengthening visual–text grounding and bilingual reasoning for downstream misinformation detection.
 Stage 2: Instruction Tuning.
@@
 We follow a unified sequence format:
-[…]
+$X=\texttt{<system>}\,[X_{\text{q}};X_{v}]\,\texttt{<STOP>}\,X_{\text{a}}\,\texttt{<STOP>}$
 where $X_{\text{q}}$ contains the natural language instruction and bilingual subtitles, and $X_{\text{a}}$ consists of three binary labels and a free-form explanation, formatted as:
-[…]
+$\texttt{<think>}\,E\,\texttt{</think>}\quad\texttt{<answer>}\,y\,\texttt{</answer>}$
 The model is trained to generate $X_{\text{a}}$ conditioned on $X_{\text{q}}$ and $X_{v}$, by maximizing the likelihood:
@@
 Only the tokens in $X_{\text{a}}$ are used to compute the loss.
-This tuning step encourages the model to align with the structure and reasoning required for fine-grained misinformation detection.
+This tuning step encourages the model to align with the structure and reasoning required for misinformation detection.
 Stage 3: GRPO-based Reasoning Optimization.
-To improve the model’s reasoning quality and explanation consistency, we apply Group Relative Policy Optimization (GRPO) [12], a reinforcement learning strategy that adjusts the policy by comparing outputs within a batch rather than relying on absolute rewards alone.
+To enhance reasoning and explanation quality for multimodal misinformation detection, we adopt Group Relative Policy Optimization (GRPO) [shao2024deepseekmath].
+GRPO ranks candidate outputs within each batch, which fits our setting where explanations for subtle visual edits and bilingual subtitle inconsistencies may be partially correct, and avoids the reward-model sensitivity of PPO and the pairwise preference assumption of DPO.
 Given a question $q$, GRPO samples $N$ candidate responses $\{r_{1},r_{2},\ldots,r_{N}\}$ from the policy $\pi_{\theta}$ and evaluates each response $o_{i}$ using a reward function $R\left(q,o_{i}\right)$, which measures the quality of the candidate in the context of the given question.
 GRPO encourages the model to generate responses with higher advantages within the group by updating the policy $\pi_{\theta}$ using the following objective:
-$\mathcal{J}_{\mathrm{GRPO}}(\theta)=\mathbb{E}\left[{\{o_{i}\}_{i=1}^{N}\sim\pi_{\theta_{\text{old}}}(q)}\right]\frac{1}{N}\sum_{i=1}^{N}\left\{\min\left[s_{1}\cdot A_{i},\;s_{2}\cdot A_{i}\right]-\beta\mathbb{D}_{\mathrm{KL}}\left(\pi_{\theta}\|\pi_{\mathrm{ref}}\right)\right\}$ (2)
-$A_{i}$ $=\frac{r_{i}-\mathrm{mean}\{r_{1},r_{2},\ldots,r_{N}\}}{\mathrm{std}\{r_{1},r_{2},\ldots,r_{N}\}},\quad s_{1}=\frac{\pi_{\theta}(o_{i}\mid q)}{\pi_{\theta_{\text{old}}}(o_{i}\mid q)},\quad s_{2}=\mathrm{clip}\left(s_{1},\;1+\epsilon,\;1-\epsilon\right),$ (3)
+$\mathcal{J}_{\mathrm{GRPO}}(\theta)$ $=\mathbb{E}_{\{o_{i}\}_{i=1}^{N}\sim\pi_{\theta_{\text{old}}}(q)}\bigg[\frac{1}{N}\sum_{i=1}^{N}\Big(\min\!\big(s_{1}A_{i},\;s_{2}A_{i}\big)$ (2)
+$\qquad-\beta\,\mathbb{D}_{\mathrm{KL}}\!\big(\pi_{\theta}\,\|\,\pi_{\mathrm{ref}}\big)\Big)\bigg].$
+$A_{i}$ $=\frac{r_{i}-\mathrm{mean}(r_{1},\ldots,r_{N})}{\mathrm{std}(r_{1},\ldots,r_{N})},$ (3)
+$s_{1}$ $=\frac{\pi_{\theta}(o_{i}\mid q)}{\pi_{\theta_{\text{old}}}(o_{i}\mid q)},\qquad s_{2}=\mathrm{clip}\!\left(s_{1},\;1-\epsilon,\;1+\epsilon\right).$
 Where $A_{i}$ represents the advantage of the candidate response $o_{i}$ relative to other sampled responses. Following DeepSeek-R1, we use both format and accuracy reward.
@@
 Format reward.
-To enforce structured outputs, we define a format reward $R_{\text{format}}$ that checks whether the model response follows a predefined format, using <answer></answer> tags for predictions and an optional <think></think> section for explanations.
-This ensures consistency across outputs and simplifies automatic evaluation.
-We use regular expressions for validation: $R_{\text{format}}=\begin{cases}0,&\text{if output matches expected format}\\ 1,&\text{otherwise}\end{cases}$.
-Localization reward.
-We use an IoU-based reward $R_{\text{Loc}}$ to assess the alignment between predicted and ground-truth tampered regions.
-To simplify computation, both are represented as bounding boxes.
-The reward is defined as: $R_{\text{Loc}}=\frac{|\mathcal{M}_{\text{pred}}\cap\mathcal{M}_{\text{gt}}|}{|\mathcal{M}_{\text{pred}}\cup\mathcal{M}_{\text{gt}}|}$, where $\mathcal{M}_{\text{pred}}$ and $\mathcal{M}_{\text{gt}}$ denote the predicted and ground-truth bounding boxes, respectively.
+To enforce structured outputs, we define a format reward $R_{\text{format}}$ that equals 1 when the model output follows the predefined format with <answer></answer> and optional <think></think> tags, and 0 otherwise: $R_{\text{format}}=\mathds{1}\!\left[\text{output matches expected format}\right].$ Localization reward.
+We define an IoU-based reward $R_{\text{Loc}}=\frac{|\mathcal{M}_{\text{pred}}\cap\mathcal{M}_{\text{gt}}|}{|\mathcal{M}_{\text{pred}}\cup\mathcal{M}_{\text{gt}}|}$, where $\mathcal{M}_{\text{pred}}$ and $\mathcal{M}_{\text{gt}}$ are the predicted and ground-truth bounding boxes of the tampered region.
 Classification reward.
-To encourage accurate decision-making in multimodal misinformation detection, we define a classification reward $R_{\text{cls}}$ that compares the model’s predicted label with the ground-truth annotation.
-This applies to tasks such as detecting whether the image is manipulated or whether the subtitles are semantically consistent.
-$R_{\text{cls}}=\begin{cases}1,&\text{if }C_{\text{pred}}=C_{\text{gt}}\\ 0,&\text{otherwise}\end{cases}$.
+We define $R_{\text{cls}}=\mathds{1}\!\left[C_{\text{pred}}=C_{\text{gt}}\right].$ to give a reward of 1 when the predicted label matches the ground truth and 0 otherwise.
 The final reward combines all task-specific objectives, including formatting, classification, localization: $R_{\text{total}}=R_{\text{format}}+R_{\text{cls}}+R_{\text{loc}}$.
-This unified reward encourages the model to generate structured, accurate, and interpretable predictions across modalities.
+This unified reward encourages the model to generate structured, accurate, and interpretable predictions across modalities without introducing any additional weighting coefficients, so each objective contributes equally.
 This progressive training strategy equips BiMi with the capability to perform fine-grained classification, localization, and explanation, and is designed to support generalization to challenging, real-world misinformation cases.
@@
 5.1 Experimental Setup
-We evaluate BiMi on BiMiBench and MMFakeBench [8].
-On BiMiBench, the model performs six-class classification, tampered region localization, and explanation generation.
-For MMFakeBench, which lacks subtitle information, we follow its four-class setting and provide text context via prompts.
+Table 3: Combined results on BiMiBench and MMFakeBench (metrics in %).
+IoU is reported when region-level localization is available (“–” denotes not applicable).
+All results are averaged over three runs with standard deviation below 0.5%.
+Bold and underline indicate best and second-best performance, respectively.
+Method MLLM BiMiBench MMFakeBench
+ACC F1 IoU BERT ACC F1
+InternVL3 [zhu2025internvl3exploringadvancedtraining] Qwen2.5-7B 20.85 (-25.95) 18.42 (-24.37) 7.23 (-23.01) 71.82 (-11.08) 42.69 (-28.25) 34.80 (-27.69)
+InternVL3 [zhu2025internvl3exploringadvancedtraining] Qwen2.5-14B 28.47 (-18.33) 26.71 (-16.08) 14.38 (-15.86) 76.54 (-6.36)
+Qwen3 [qwen3] Qwen3-8B 21.43 (-25.37) 14.96 (-27.83) 6.28 (-23.96) 73.92 (-8.98) 44.82 (-26.12) 39.49 (-23.00)
+LLaVA-1.6 [liu2024llavanext] Vicuna-7B 22.91 (-23.89) 20.74 (-22.05) 8.91 (-21.33) 75.77 (-7.13) 30.09 (-40.85) 24.31 (-38.18)
+Llama 3.1 [grattafiori2024llama3herdmodels] Llama-3.1-8B 24.61 (-22.19) 21.94 (-20.85) 5.27 (-25.0) 72.91 (-9.99) 32.67 (-38.27) 28.30 (-34.19)
+Gemma3 [team2025gemma] Gemma3-4B-IT 17.48 (-29.32) 11.73 (-31.06) 4.29 (-25.95) 70.84 (-12.06) 27.19 (-43.75) 22.04 (-40.45)
+SNIFFER [qi2024sniffer] Vicuna-13B 37.90 (-8.90) 33.63 (-9.16) 80.41 (-2.49) 67.49 (-3.45) 61.44 (-1.05)
+MMD-Agent [liu2024mmfakebench] Qwen2.5-7B 34.78 (-12.02) 29.48 (-13.31) 62.78 (-8.16) 52.83 (-9.66)
+BiMi (Ours) Gemma3-4B-IT 46.80 42.79 30.24 82.90 70.94 62.49
+We evaluate BiMi on BiMiBench and MMFakeBench [liu2024mmfakebench].
+On BiMiBench, the model performs six-class misinformation classification, region-level tamper localization, and explanation generation.
+For MMFakeBench, which does not provide subtitles, we follow its standard four-class setting and supply minimal textual context via a unified prompt; subtitle-specific objectives are disabled while the rest of the pipeline remains unchanged.
 Implementation Details.
-We use the Gemma 3 as the LLM, applying our three-stage post-training pipeline.
-All experiments are run on 48GB V100 GPUs with a batch size of 16, learning rate $1\times 10^{-5}$, and frozen vision encoder.
-GRPO is applied for two additional epochs.
-We design structured instruction prompts to guide the model in tasks, as detailed in Table 3.
-Please analyze the given image containing both Chinese and English subtitles and complete the following three tasks: (1) Classification Task: classify the alignment between the image and the subtitles into one of the following six categories: "all consistent", "image manipulated", "both subtitles misaligned with image", "only English aligned", "only Chinese aligned", "all inconsistent".
-(2) Manipulation Detection: if the image has been manipulated, return one or more bounding boxes for the manipulated regions in the format: {"bbox":[x_min, y_min, x_max, y_max]}.
-If no manipulation is found, return an empty list.
-(3) Decision Explanation: briefly explain your thinking before the classification and any detected regions.
-Return your output using the following format, wrapped in tags:
-<think>Your explanation here.</think>
-<answer>{"classification": "result", "region": [{"bbox": [x_min, y_min, x_max, y_max]}]}</answer>
-Table 3: Prompt templates used in BiMi for multimodal misinformation detection
+We adopt Gemma-3 [team2025gemma] as the base multimodal large language model and apply the proposed three-stage post-training pipeline: domain align, supervised fine-tuning (SFT) and GRPO-based reinforcement learning.
+The vision encoder is kept frozen throughout post-training, while only the multimodal adapter and language-model parameters are updated.
+All experiments are conducted on 48 GB NVIDIA V100 GPUs with a global batch size of 16 and an initial learning rate of $1\times 10^{-5}$.
+SFT is run for 3 epochs with early stopping based on validation loss, followed by 2 additional epochs of GRPO training.
+We fix the random seed to 42 and use AdamW with a weight decay of $1\times 10^{-2}$.
+The final reward combines the format, classification, and localization components without introducing any weighting coefficients, i.e., $R_{\text{total}}=R_{\text{format}}+R_{\text{cls}}+R_{\text{loc}}$, so that each objective contributes equally.
 Baselines.
-We compare BiMi against a range of strong multimodal baselines, including InternVL3 [27], Qwen3 [47], LLaVA-1.6 [33], and LLama3.1 [48], which represent leading approaches in vision-language modeling.
-We also include specialized misinformation detection systems such as SNIFFER [7] and MMD-Agent [8], designed for out-of-context and multimodal misinformation.
+We compare BiMi against a range of strong multimodal baselines, including InternVL3 [zhu2025internvl3exploringadvancedtraining], Qwen3 [qwen3], LLaVA-1.6 [liu2024llavanext], and LLama3.1 [grattafiori2024llama3herdmodels], which represent leading approaches in vision-language modeling.
+We also include misinformation detection systems such as SNIFFER [qi2024sniffer] and MMD-Agent [liu2024mmfakebench], designed for out-of-context and multimodal misinformation.
 All models are evaluated using the same inputs: the image with overlaid Chinese-English subtitles and a task-specific prompt.
-For ablation, we test BiMi variants without GRPO, without retrieval, and with alternative MLLM backbones.
 Evaluation Metrics.
-We report Accuracy (ACC) and F1 score for classification, IoU for localization.
-Inspired by [30], we evaluate explanation quality via BERTScore [49] between response and reference explanations generated by prompting GPT-4o [50] with labels and manually verified for correctness.
-Table 4: Main results on BiMiBench. BiMi achieves the best performance across all metrics. Bold and underline denote best and second-best results.
-Method LLM Res. ACC F1 IoU BERT
-InternVL3 [27] Qwen2.5-7B 224 20.85 18.42 7.23 71.82
-InternVL3 [27] Qwen2.5-14B 224 28.47 26.71 14.38 76.54
-Qwen3 [47] Qwen3-8B 224 21.43 14.96 6.28 73.92
-LLaVA-1.6 [33] Vicuna-7B 224 22.91 20.74 8.91 75.77
-LLama3.1 [48] Llama-3.1-8B 224 24.61 21.94 5.27 72.91
-Gemma3 [48] Gemma3-4B-IT 336 17.48 11.73 4.29 70.84
-SNIFFER [7] Vicuna-13B 224 37.90 33.63 80.41
-MMD-Agent [8] Qwen2.5-7B 224 34.78 29.48
-BiMi Gemma3-4B-IT 336 46.80 42.79 30.24 82.90
+We report Accuracy (ACC) and F1 for classification and IoU for tampered-region localization.
+For explanation quality, we follow Liu et al.
+[liu2023llava] and compute BERTScore [bert-score] between model outputs and pseudo-references generated by GPT-4o [hurst2024gpt].
+Pseudo-references are produced via structured prompting with label-grounded templates and manually verified by five trained annotators to ensure factual correctness and strict alignment with the ground-truth manipulations.
+To confirm that BERTScore reflects human judgment, the same annotators also rated 400 randomly sampled explanations on a 5-point Likert scale across five ground-truth–aligned dimensions—subtitle alignment, visual detail consistency, reasoning consistency, clarity and readability, and completeness.
+The high agreement between these human ratings (Fleiss’ $\kappa=0.71$) demonstrates the reliability and practical relevance of our evaluation protocol.
+More evaluation details are provided in the Appendix.
 5.2 Main Results
-Table 5: MMFB results.
-Method ACC F1
-InternVL3 42.69 34.80
-Qwen3 44.82 39.49
-LLaVA-1.6 30.09 24.31
-LLama3.1 32.67 28.30
-Gemma3 27.19 22.04
-SNIFFER 67.49 61.44
-MMD-Agent 62.78 52.83
-BiMi 70.94 62.49
 Performance Comparison.
-We compare BiMi with strong baselines across three subtasks on the BiMiBench test set: (1) misinformation detection, (2) tampered region localization, and (3) explanation generation.
-Results are summarized in Table 4.
-BiMi significantly outperforms all baselines in both all tasks.
-In particular, it achieves a +8.9 accuracy gain in image-subtitle consistency, a +15.9 localization accuracy, and a +2.5 BERTScore improvement in explanation generation over the strongest baseline.
-BiMi’s performance also benefits from the GRPO reward design, which improves explanation specificity and localization precision.
-The results on MMFakeBench (4-class classification task) are shown in Table 5.
+We evaluate BiMi across three key subtasks on BiMiBench—(1) misinformation detection, (2) tampered region localization, and (3) explanation generation—and additionally assess cross-dataset generalization on MMFakeBench.
+As summarized in Table 3, BiMi consistently outperforms all baseline MLLMs across every metric and benchmark.
+On BiMiBench, it achieves notable improvements over the strongest baseline, including a +8.9 gain in detection accuracy, a +15.9 improvement in localization IoU, and a +2.5 increase in BERTScore for explanation quality.
+These gains are further strengthened by our GRPO-based reward design, which enhances both localization precision and explanation specificity.
+BiMi also achieves the best performance on MMFakeBench (4-class classification task), demonstrating strong robustness and generalization beyond the training domain.
+Computation and Efficiency.
+End-to-end latency includes three stages: model forward pass, OCR subtitle extraction, and external retrieval.
+Pure model inference averages  320 ms per sample, while enabling OCR and web search adds 0.8–1.2 s depending on network conditions—similar to typical retrieval-augmented MLLMs.
+To ensure robustness, BiMi adopts a graceful-degradation strategy: if OCR or retrieval times out, the system automatically falls back to internal reasoning with minimal performance impact.
 Ablation Study.
@@
 (ii) Instruction Tuning.
-Eliminating instruction tuning leads to a drastic performance drop, accuracy falls by 31.18 and BERTScore by 12.06, highlighting its importance in enabling structured bilingual understanding.
+Eliminating instruction tuning leads to a drastic performance drop, accuracy falls by 31.18 and BERTScore by 12.06, highlighting its importance in enabling bilingual understanding.
 (iii) GRPO.
@@
 (iv) Output Order.
-Reversing the output order (classify before explain) slightly lowers accuracy (–1.94) and BERTScore (–1.11), suggesting that generating explanations first facilitates more coherent reasoning.
+Reversing the output order (classify before explain) slightly lowers accuracy (–1.94) and BERTScore (–1.11), suggesting that generating explanations facilitates more reasoning.
 Table 4 summarizes the results.
 These findings confirm that all components are essential, with instruction tuning and GRPO being particularly critical for boosting accuracy and explanation quality.
+Figure 5: Comparison of explanations from InternVL3 (middle) and our model (right). Top left: input; bottom left: original sample. Some responses are truncated due to space constraints.
 5.3 Explainability Analysis
 Beyond accurate classification and localization, generating high-quality natural language explanations is crucial for interpretable misinformation detection.
-We analyze explanation ability through (1) qualitative comparisons, (2) the influence of the retrieval module, and (3) evaluation under bilingual.
+We analyze explanation ability through qualitative comparisons, the influence of the retrieval module, and evaluation under bilingual.
 Explanation Quality.
-To understand BiMi’s reasoning ability, we compare its output with InternVL3 [27] on a representative case (Figure 5).
-Both models recognize that the subtitles are semantically aligned, showing bilingual understanding.
-However, InternVL3 relies on external factual cues and overlooks the mismatch between image and text.
-In contrast, BiMi grounds its decision in the image, correctly identifies the manipulation, and localizes the tampered region.
-It also provides a clearer, step-by-step explanation, showing stronger visual grounding and more faithful reasoning.
-Figure 5: Comparison of explanations from InternVL3 (middle) and our model (right).
-The top left is input.
-The bottom left is the original sample.
-Note: Due to space constraints, some model responses are truncated.
-Full examples and additional results are provided in the Appendix.
-Effect of Retrieval.
-We analyze a real-world social media post with subtle bilingual inconsistency (Figure 6).
-The removed "student" in the Chinese translation alters the meaning.
-Without retrieval, the model misclassifies due to a lack of context.
-The retrieval result provides a key clue for classification.
-BiMi correctly identifies the inconsistency, showing its benefit for nuanced cross-lingual reasoning.
+We compare BiMi with InternVL3 [zhu2025internvl3exploringadvancedtraining] on a representative case (Figure 5).
+Both models capture the bilingual subtitle alignment, but InternVL3 relies on external cues and misses the image–text mismatch.
+BiMi instead its decision in the image, correctly localizes the manipulation, and offers a cleare explanation.
+Retrieval Module on Real Samples.
+We further tested the retrieval component on a 100-sample set of authentic bilingual news images to examine its behavior in real-world scenarios.
+Human evaluation shows that retrieved evidence improves explanation clarity or correctly resolves subtle cross-lingual inconsistencies in 9% of cases, remains neutral in 89%, and results in off-topic passages in only 2%, typically due to OCR noise in Chinese subtitles.
+Importantly, even in these rare failure cases, the final predictions remain stable, indicating that retrieval errors do not propagate into the core reasoning pipeline.
+Overall, these results suggest that the retrieval module provides timely factual context that can meaningfully enrich the model’s explanations, while its downside risk is minimal—an encouraging property for practical deployment in noisy real-world environments.
 Table 5: Accuracy and BERTScore for each language variant (4-category classification).
@@
 Bilingual multimodal capability.
-To evaluate bilingual understanding, we construct a variant of the test set where the model receives only one subtitle, Chinese or English.
-This reduces the task to a 4-category classification.
-By ignoring cross-lingual consistency, the single-language setting reduces reasoning difficulty.
-As shown in Table 7, the model performs better with English-only inputs, likely due to the predominance of English data during training.
-This highlights a gap in bilingual support and the need for stronger Chinese-language modeling in this domain.
-Figure 6: Social media example with subtle bilingual inconsistency. The Chinese translation omits "student", changing the meaning and potentially misleading Chinese readers.
+To assess bilingual understanding, we also create a single-language test variant where the model receives only Chinese or only English subtitles, reducing the task to a 4-class classification.
+Removing cross-lingual comparison makes the task substantially easier.
+As shown in Table 5, performance is noticeably higher with English-only inputs, reflecting the dominance of English data and reasoning patterns in pre-training.
+Chinese-only inputs introduce greater variance, particularly with idioms or long, multi-clause subtitles.
+These results reveal a clear gap in bilingual robustness and underscore the need for stronger Chinese modeling, improved tokenization, and more balanced multilingual data in future systems.
 Failure Cases.
-We observe that BiMi occasionally fails when manipulations are extremely subtle or when OCR fails to extract accurate subtitles.
-In such cases, either the localization is off-target, or the explanation is overly general.
-More failure cases and analysis are provided in the Appendix.
+BiMi’s remaining errors stem mainly from two sources.
+First, extremely subtle manipulations can lead to off-target localization or generic explanations, especially when visual edits or cross-lingual inconsistencies are barely perceptible.
+Second, OCR quality has a substantial impact: about 37% of misclassifications result from truncated characters or rare-character errors in Chinese subtitles, issues rarely seen in English text.
+As shown in Table 6, performance drops noticeably when using OCR-extracted subtitles instead of ground-truth text, confirming OCR noise as a major contributor.
+More examples are in the Appendix.
+Table 6: Classification accuracy (ACC) and BERTScore (%) using Ground-truth subtitles (manually provided) versus OCR-extracted subtitles (automatically recognized).
+Language Subtitle Source ACC (%) BERTScore
+English Ground-truth 84.24 88.10
+English OCR-extracted 83.02 85.92
+Chinese Ground-truth 80.07 86.54
+Chinese OCR-extracted 75.33 80.21
 6 Conclusion and Limitation
 Conclusion.
-We study the task of detecting multimodal inconsistencies in news images with bilingual (Chinese-English) subtitles.
+We study the task of detecting multimodal inconsistencies in news images with bilingual subtitles.
 To support this, we introduce BiMiBench, a large-scale benchmark with 104K samples featuring realistic visual edits and cross-lingual mismatches.
 We propose BiMi, a bilingual multimodal framework combining a three-stage post-training strategy and retrieval-augmented reasoning.
-Experiments show that BiMi achieves state-of-the-art performance in classification, localization, and explanation, highlighting its effectiveness in real-world scenarios.
-Limitation and future work.
+Experiments show that BiMi achieves state-of-the-art performance across tasks, highlighting its effectiveness in real-world scenarios.
+This work advances research on bilingual and multimodal misinformation detection.
+Limitation.
 While BiMi achieves strong results on bilingual multimodal detection, several limitations remain.
-First, the current setup focuses on Chinese-English subtitles; extending to more languages would enhance generalization.
-Second, localization is limited to bounding boxes, which may miss fine-grained manipulations—future work could explore pixel-level segmentation.
-Third, BiMiBench samples use fixed subtitle styles and editing templates; expanding visual diversity and subtitle formats would better reflect real-world cases.
-Finally, while BiMi offers solid performance, its reasoning and explanation ability can be further improved, especially in ambiguous scenarios.
-Appendix
-Section A provides details on the dataset, including: dataset statistics, misinformation generation strategy, example visualization, and manual verification.
-Section B outlines implementation and experimental details, including: model architecture and training settings, hyperparameters.
-Section C discusses the broader impact and ethical implications of our work.
-Section D presents qualitative examples.
-Section E analyzes the failure case.
-Appendix A Dataset Details
-Dataset Statistics.
+First, it currently targets only Chinese–English subtitles; extending to additional languages would improve cross-lingual generalization.
+Second, localization is restricted to bounding boxes and may miss fine-grained manipulations.
+Third, the data diversity of BiMiBench is limited: beyond broader subtitle templates and editing styles, richer sources spanning multiple languages and topical domains are needed to better reflect real-world variability.
+Fourth, retrieval introduces extra latency and depends on external sources, which can occasionally return noisy or outdated information.
+Finally, BiMi’s reasoning capability could be strengthened, especially for cases with conflicting multimodal evidence.
+Overview of Supplementary Material
+Section A: Dataset and ethics.
+Section B: Additional experiments and cases.
+Section C: Implementation details.
+Appendix A Dataset Construction and Quality Control
+A.1 Dataset Statistics
 BiMiBench comprises 104,000 samples spanning six fine-grained misinformation categories.
-Figure 7 summarizes key dataset properties.
-The left panel shows the distribution of samples across categories, ensuring diverse coverage of manipulation types.
+Figure 6 summarizes the key dataset properties.
+Beyond the balanced distribution across categories (left), BiMiBench exhibits substantial diversity in both visual content and bilingual subtitle structures.
+The images cover a wide range of news domains, including politics, disasters, sports, daily life, technology, and international affairs, ensuring that manipulation patterns are not confined to a narrow set of scenes.
+The subtitles vary in length, writing style, and translation quality across topics in both Chinese and English.
+Such multimodal richness helps simulate real-world bilingual misinformation scenarios and provides a challenging, comprehensive testbed for fine-grained manipulation detection.
 Figure 6: Dataset statistics of BiMiBench.
@@
 Right: normalized area of manipulated regions (bounding boxes height).
-Misinformation Generation Strategy.
+A.2 Misinformation Generation Strategy.
 To simulate realistic multimodal misinformation, we generate manipulated subtitles by introducing subtle yet misleading semantic alterations.
@@
 We design task-specific prompts to guide misinformation generation, ensuring fluency and semantic plausibility while introducing factual inconsistencies; an example prompt used in this process is shown in Table 7.
-<system> Based on the image, create a fictional, sensational news headline or summary in one sentence.
+Table 7: Prompt templates used in BiMi for multimodal misinformation detection
+<<system>> Based on the image, create a fictional, sensational news headline or summary in one sentence.
 The story must be clearly fake but sound plausible, and must be no more than 30 words.
 The news should sound outrageous, scandalous, or unbelievable, yet still written in a serious, news-like tone.
-</system>
-Table 8: Prompt templates used in BiMi for multimodal misinformation detection
-Example Visualization. Figure 8 illustrates representative BiMiBench samples, covering various manipulation types, language inconsistencies, and corresponding ground truth labels.
-Figure 8: Sample BiMiBench examples with bilingual subtitles and visual or text manipulations.
-Manual Verification.
+<</system>>
+Data Diversity.
+To ensure broad and representative diversity in the synthetic misinformation, we introduce variation at three complementary stages.
+Data source: we begin with the large and professionally curated VisualNews corpus, whose wide topical coverage—politics, economics, sports, culture—naturally provides a variety of writing styles and visual contexts, ensuring that manipulations are not limited to a narrow set of news domains.
+Editing: on top of this diverse base we generate multiple kinds of factual distortions, ranging from single–word entity swaps to complex changes in temporal references, numerical values and causal relations.
+We also randomize prompt formulations and the sampling parameters of the generation model so that the manipulated bilingual subtitles exhibit heterogeneous and unpredictable inconsistencies.
+Label assignment: after generation, each candidate sample is manually reviewed by trained annotators and assigned to one of the six defined misinformation categories, reflecting the precise combination of visual edits and cross-lingual subtitle inconsistencies.
+This three-stage process guarantees that BiMiBench covers a wide spectrum of realistic manipulation patterns and prevents overfitting to any type of false information.
+A.3 Review Protocol
 To ensure data quality and annotation reliability, all samples in BiMiBench were manually verified by a team of five trained annotators, each with backgrounds in journalism, linguistics, or media studies, and prior experience with fact-checking or misinformation analysis.
@@
 This multi-layered, expert-driven verification process ensures that BiMiBench maintains high annotation quality, semantic precision, and supports reliable benchmarking for multilingual multimodal misinformation detection.
-Appendix B Implementation Details
+A.4 Ethical and Licensing Considerations
+BiMiBench is derived from the publicly available VisualNews corpus, which provides professionally curated image–text news pairs under a license permitting academic research.
+All original images and captions remain the intellectual property of their respective news organizations; our release is strictly for non-commercial, research purposes and requires users to acknowledge the original sources and this benchmark in any derivative work.
+While BiMiBench is designed to advance scientific understanding of multimodal misinformation detection, it necessarily contains intentionally manipulated content: images edited to introduce subtle visual changes and bilingual subtitles rewritten to create semantic inconsistencies.
+This “dual nature”—a resource built for research, yet embedding realistic examples of fabricated content—poses an inherent risk of misuse if taken out of context or circulated without clear academic framing.
+To mitigate these risks, we (i) clearly mark all manipulated samples as synthetic, (ii) provide the data only for research and educational use, and (iii) require that any redistribution or downstream work include appropriate citations and comply with relevant copyright and data-protection regulations.
+We explicitly prohibit using BiMiBench to create or disseminate deceptive media outside controlled research settings.
+These safeguards aim to ensure that the benchmark remains a tool for studying and countering misinformation rather than a source of it.
+A.5 Example Visualization.
+Figure 7 illustrates representative BiMiBench samples, showcasing a wide range of manipulation patterns, cross-lingual inconsistencies, and subtitle complexities.
+Figure 7: Sample BiMiBench examples with bilingual subtitles and visual or text manipulations.
+Appendix B Additional Experiments
+B.1 Reference Generation
+To construct high-quality reference explanations for automatic evaluation, we adopt a two-stage human–LLM pipeline.
+First, for each sample we prompt GPT-4o with task-specific, label-grounded templates to generate two candidate explanations conditioned on the manipulated modality and misinformation type.
+Second, a team of five trained annotators—each with a background in journalism, media studies, or NLP—independently reviewed the two candidates and selected the one that best satisfied five ground-truth–aligned dimensions: (i) Subtitle alignment, (ii) Visual detail consistency, (iii) Reasoning consistency, (iv) Clarity and readability, and (v) Completeness.
+The chosen explanation for each sample was retained as the pseudo-ground truth used in BERTScore evaluation.
+This procedure ensures that every reference explanation is both fluent and factually faithful while keeping the generation process efficient and reproducible.
+Table 8: Human evaluation criteria and scoring scheme for selecting reference explanations.
+Annotators scored each criterion on a 0–2 scale (0 = not satisfied, 1 = partially satisfied, 2 = fully satisfied), and explanations with a total score below 7 were discarded or revised.
+Criterion Description Score
+Subtitle alignment Checks consistency between English–Chinese subtitles and the described manipulation. 0–2
+Visual detail consistency Ensures the explanation correctly reflects visible tampered regions or edits. 0–2
+Reasoning consistency Evaluates logical coherence between detected evidence and the explanation. 0–2
+Clarity and readability Measures fluency and conciseness suitable for news-style explanations. 0–2
+Completeness Ensures all manipulated elements are covered without omitting key evidence. 0–2
+B.2 Performance on Real-World Samples
+To assess real-world generalization, we manually collected 100 recent social-media posts spanning political, health, and international news.
+These posts contain image–caption manipulations, cross-lingual inconsistencies, and subtle visual edits typical of in-the-wild misinformation.
+Because the authenticity of such posts cannot be verified with ground-truth labels, the 100 examples were expert-curated: five annotators with backgrounds in journalism and fact-checking jointly reviewed the content and reached consensus on whether each case constitutes misinformation.
+This human consensus served as the reference label set.
+Evaluated against these expert judgments, BiMi correctly identified 81 of the 100 curated cases (81% accuracy) .
+These findings suggest that BiMi can transfer effectively from synthetic benchmarks to uncontrolled real-world conditions, even when the ground truth is based on expert assessment rather than definitive factual verification.
+We analyze a real-world social media post with subtle bilingual inconsistency (Figure 8).
+The removed ”student” in the Chinese translation alters the meaning.
+Without retrieval, the model misclassifies due to a lack of context.
+The retrieval result provides a key clue for classification.
+BiMi correctly identifies the inconsistency, showing its benefit for nuanced cross-lingual reasoning.
+Figure 8: Social media example with subtle bilingual inconsistency.
+The Chinese translation omits ”student”, changing the meaning and potentially misleading Chinese readers.
+Middle: InternVL3 result; Right: BiMi result.
+B.3 Effect of OCR Module
+This section provides brief supplementary observations expanding on the OCR analysis in the main paper.
+While the core quantitative results are already reported, here we include additional qualitative notes.
+During manual inspection, we found that OCR errors most commonly occur in Chinese subtitles containing long sentence structures or low-resolution text, which aligns with the degradation patterns shown in the main paper.
+In contrast, English subtitles are generally recognized cleanly, leading to smaller performance differences between ground-truth and OCR-extracted inputs.
+In many real-world samples, OCR inaccuracies do not alter the final classification category but may affect explanation specificity.
+We therefore recommend simple preprocessing steps—such as cropping subtitle regions or modest contrast enhancement—to stabilize OCR quality when applying the model outside the benchmark setting.
+These findings show that BiMi remains largely robust to moderate OCR noise; the observed degradation originates mainly from preprocessing rather than from limitations in the model’s bilingual reasoning.
+When clean Chinese subtitles are provided, BiMi’s accuracy approaches the English benchmark.
+Improving OCR fidelity can enhance overall system reliability in real-world deployments.
+B.4 Analysis of Bilingual Performance Gap
+BiMi achieves higher accuracy on English-only inputs (82.48% ACC) than on Chinese-only inputs (72.32% ACC).
+To disentangle the causes of this gap, we conducted a controlled evaluation using ground-truth subtitles versus OCR-extracted subtitles, as summarized in Table 9.
+Table 9: Impact of OCR on classification accuracy.
+Language GT (%) OCR (%) Drop (%)
+English 84.2 83.0 -1.2
+Chinese 80.0 75.3 -4.7
+The results show that Chinese performance is far more sensitive to OCR noise, with a 4.7% drop compared to only 1.2% for English.
+Thus, OCR quality in Chinese subtitles is the primary factor behind the overall 10% gap.
+Secondary contributors include (i) training data imbalance: VisualNews contains more English captions, which biases the backbone model toward English, and (ii) language-model priors: although Gemma 3 supports bilingual reasoning, its representations are stronger for English.
+When clean Chinese subtitles are provided, BiMi’s accuracy approaches the English benchmark, confirming that OCR errors are the dominant bottleneck.
+To close this gap, future work will focus on (1) enhancing Chinese OCR through confidence-based filtering and multi-OCR consensus, (2) exploring LLM-based subtitle repair to recover noisy outputs, and (3) fine-tuning on more balanced bilingual data.
+Overall, the performance difference is mainly driven by OCR noise, with data imbalance and model priors as secondary factors.
+B.5 Computation and Efficiency
+BiMi employs a lightweight multimodal backbone (Gemma 3–4B) to maintain manageable training and inference costs.
+All experiments are conducted on a single NVIDIA V100 48 GB GPU.
+End-to-end inference latency is determined by three components: the model forward pass, OCR-based subtitle extraction, and the external retrieval module.
+Without retrieval, pure model inference averages 320 ms per sample.
+Enabling the full retrieval pipeline—including OCR and Google Search API queries—adds approximately 0.8–1.2 s of overhead, depending on network conditions and hardware.
+This additional cost remains acceptable for offline evaluation and is broadly consistent with other retrieval-augmented MLLMs.
+To avoid retrieval becoming a single point of failure, BiMi incorporates a graceful-degradation mechanism: if OCR or web search times out or returns no relevant context, the model skips the retrieved evidence and falls back to pure visual–textual reasoning, keeping classification and explanation performance stable.
+During training, the vision encoder is frozen and only LoRA adapters are updated, which keeps GPU memory usage below 40 GB and allows a batch size of 16.
+These design choices ensure that BiMi can be reproduced on standard academic hardware while maintaining reasonable inference latency and memory footprint.
+B.6 Failure Case Study
+Despite the overall effectiveness of our model, we observed several notable failure cases that highlight its current limitations.
+One recurring issue involves semantic misinterpretation, particularly when the input contains rare or ambiguous phrases.
+For example, in a prompt requiring multi-step reasoning or contextual understanding (e.g., “Describe the implication of climate change on coastal microeconomies”), the model occasionally generates generic or overly broad responses that fail to capture the specific nuances of the query.
+Another type of failure arises in edge cases involving domain-specific knowledge.
+Figure 9 shows a failure case where the model misclassified an image manipulation example as “All Inconsistent.” The English caption aligns with the image, but the Chinese one introduces a false espionage narrative.
+Key issues include (1) misclassification, (2) incorrect caption judgment (only the Chinese caption was wrong), and (3) a missing bounding box.
+Figure 9: Failure case
+B.7 Qualitative Cases
+Figures 10, 11, 12, 13, 14, 15, 16, 17 show the examples.
+Appendix C Implementation Details
+C.1 Training
 Model Architecture.
 Our framework is built upon the publicly released Gemma 3 multimodal language model.
-The vision encoder is initialized from a pretrained ViT-L model, and remains frozen throughout all training stages.
-LoRA parameters are fine-tuned.
-The model accepts the inputs of the image with English and Chinese subtitle, which serialized into a unified prompt format.
-The image is encoded through a unified visual-language backbone.
+The vision encoder is initialized from a pretrained ViT-L and kept frozen in all training stages, while LoRA adapters are fine-tuned on top of the language backbone to reduce computational cost.
+The model accepts as input the news image with its English–Chinese subtitles, serialized into a unified prompt that embeds the subtitles as text tokens and the image as patch-level embeddings.
+Visual tokens from the frozen encoder are projected into the language model’s embedding space and fused with textual tokens through the cross-modal attention layers of Gemma 3, enabling joint reasoning over visual regions and bilingual text.
+This design allows efficient adaptation to our task while preserving the strong multilingual and vision–language alignment of the pretrained model.
 Tokenization and Input Format.
-Both English and Chinese subtitles are tokenized using the SentencePiece tokenizer aligned with the Gemma 3 backbone.
-The input is formatted into structured prompts that specify the prediction task, marking different modalities.
-Input length is capped at 512 tokens.
-Hyperparameter Summary.
-We fine-tune BiMi using the Gemma 3 model (4B) with a batch size of 8, a learning rate of $1\times 10^{-5}$, and a maximum input length of 512 tokens.
-The model is optimized with AdamW, using $\beta_{1}=0.9$, $\beta_{2}=0.98$, and a weight decay of 0.01.
-To stabilize training, we use gradient accumulation (4 steps) and mixed precision (FP16).
-The GRPO reward coefficient $\lambda$s are all set to 1.
-These hyperparameters are kept consistent across all three training stages unless otherwise specified.
+Both English and Chinese subtitles are tokenized using the SentencePiece tokenizer aligned with the Gemma 3 backbone to ensure consistent multilingual encoding.
+The input is formatted into a structured prompt that specifies the prediction task and clearly marks different modalities; visual features from the frozen ViT-L encoder are projected into the same embedding space and appended as a sequence of visual tokens.
+The combined text and image tokens are then fed to the cross-modal transformer, with the total input length (text + visual tokens) capped at 512 text tokens plus the visual tokens.
+[…]
 Prompt Template.
@@
 This format ensures consistent output structure, supporting multimodal reasoning (via <think>), interpretable classification, and region-level manipulation localization.
-<system> You are an expert in misinformation detection area.
+C.2 Retrieve Module
+To improve adaptability to rapidly emerging misinformation, we design a retrieval module that operates only at inference time and does not require additional training.
+First, we apply OCR to extract both the Chinese and English subtitles embedded in the news image.
+The two subtitles are concatenated to form a bilingual query, which is then sent to the Google Search API.
+The API returns up to the top-3 relevant passages; if no relevant documents are found, the retrieval component simply returns an empty string.
+The retrieved text snippets are concatenated and prepended to the model’s input prompt, yielding a unified representation $\mathcal{P}=\operatorname{concat}(R,I,S)$, where $R$, II and $S$ denote the retrieved context, the image and the bilingual subtitles, respectively.
+This augmented prompt is fed to the MLLM backbone so that cross-attention layers can jointly encode the external evidence and the visual–textual cues.
+Because the module only provides auxiliary context and does not participate in training, the final predictions remain primarily grounded in the image–subtitle content while benefiting from timely external information when available.
+Table 10: Prompt templates used in BiMi for multimodal misinformation detection
+<<system>> You are an expert in misinformation detection area.
 A conversation between User and Assistant.
@@
 The assistant first thinks about the reasoning process in the mind and then provides the user with the answer.
-The reasoning process and answer are enclosed within <think> </think> and <answer> </answer> tags, respectively, i.e., <think> reasoning process here </think><answer> answer here </answer>.
+The reasoning process and answer are enclosed within <<think>> <</think>> and <<answer>> <</answer>> tags, respectively, i.e., <<think>> reasoning process here <</think>><<answer>> answer here <</answer>>.
 Assign one of the following six categories based on their mutual alignment: All Consistent: The image aligns well with both Chinese and English captions.
@@
 All Inconsistent: The image, Chinese caption, and English caption are all manipulated and mutually inconsistent.
-</system> <user> Please analyze the given image containing both Chinese and English subtitles and complete the following three tasks: (1) Classification Task: classify the alignment between the image and the subtitles into one of the following six categories: "all consistent", "image manipulated", "both subtitles misaligned with image", "only English aligned", "only Chinese aligned", "all inconsistent".
+<</system>> <<user>> Please analyze the given image containing both Chinese and English subtitles and complete the following three tasks: (1) Classification Task: classify the alignment between the image and the subtitles into one of the following six categories: "all consistent", "image manipulated", "both subtitles misaligned with image", "only English aligned", "only Chinese aligned", "all inconsistent".
 (2) Manipulation Detection: if the image has been manipulated, return one or more bounding boxes for the manipulated regions in the format: {"bbox":[x_min, y_min, x_max, y_max]}.
@@
 (3) Decision Explanation: briefly explain your thinking before the classification and any detected regions.
-Return your output using the following format, wrapped in tags: <think> Your explanation here.
-</think> <answer> "classification": "result", "region": [{"bbox": [x_min, y_min, x_max, y_max]}] </answer> </user>
-Table 9: Prompt templates used in BiMi for multimodal misinformation detection
-Appendix C Broader Impact and Ethics
-BiMi is designed to enhance the interpretability and reliability of multimodal misinformation detection, particularly in the context of bilingual news media.
-By providing both classification outputs and natural language explanations, BiMi promotes transparency in AI decision-making and supports fact-checkers, journalists, and platform moderators in identifying subtle manipulations across images and subtitles.
-To ensure ethical standards, all data samples in BiMiBench have been manually reviewed, linguistic clarity, and annotation consistency.
-Low-quality, ambiguous content has been excluded.
-The dataset construction avoids personally identifiable information and sensitive content.
-BiMi is detection-focused and does not produce or promote fake content, reducing the risk of misuse.
-Nonetheless, we recognize that any publicly released model or dataset could potentially be repurposed.
-To mitigate this, we include usage guidelines and encourage responsible deployment, especially in downstream applications that may involve high-stakes decisions.
-We believe this work contributes positively by advancing research on robust, interpretable AI tools for combating harmful information while maintaining strong ethical safeguards throughout the development and release process.
-Appendix D Qualitative Examples
-Figure 9 10 11 12 13 14 15 16 show the examples.
-Figure 9: Case of All Consistency.
+Return your output using the following format, wrapped in tags: <<think>> Your explanation here.
+<</think>> <<answer>> "classification": "result", "region": [{"bbox": [x_min, y_min, x_max, y_max]}] <</answer>> <</user>>
+Figure 10: Case of All Consistent.
 Figure 11: Case of Chinese Misaligned.
-Figure 11: Case of Image Manipulation.
+Figure 12: Case of Image Manipulated.
 Figure 13: Case of English Misaligned.
@@
 Figure 17: Case of Both Misaligned.
-Appendix E Failure Cases
-Despite the overall effectiveness of our model, we observed several notable failure cases that highlight its current limitations.
-One recurring issue involves semantic misinterpretation, particularly when the input contains rare or ambiguous phrases.
-For example, in a prompt requiring multi-step reasoning or contextual understanding (e.g., “Describe the implication of climate change on coastal microeconomies”), the model occasionally generates generic or overly broad responses that fail to capture the specific nuances of the query.
-Another type of failure arises in edge cases involving domain-specific knowledge.
-Figure 17 shows a failure case where the model misclassified an image manipulation example as “All Inconsistent.” The English caption aligns with the image, but the Chinese one introduces a false espionage narrative.
-Key issues include: misclassification; incorrect caption judgment—only the Chinese caption was wrong; missing bounding box.
-This highlights weaknesses in manipulation detection, multilingual judgment, and grounding.
-Figure 17: Failure case